### v1

# Asymptotic geometry at infinity of quiver varieties

Abstract.
Using an approach developed by Melrose to study the geometry at infinity of the Nakajima metric on the reduced Hilbert scheme of points on $\mathbb{C}^{2}$, we show that the Nakajima metric on a quiver variety is quasi-asymptotically conical ($\operatorname{QAC}$) whenever its defining parameters satisfy an appropriate genericity assumption. As such, it is of bounded geometry and of maximal volume growth. Being $\operatorname{QAC}$ is one of two main ingredients allowing us to use the work of Kottke and the second author to compute its reduced $L^{2}$-cohomology and prove the Vafa-Witten conjecture. The other is a spectral gap for the Hodge-deRham operator associated with an exact $3$-Sasakian wedge metric that we obtain using the Weitzenböck formula of Semmelmann and Weingart on quaternionic-Kähler manifolds. This yields in particular a generalization and a different proof of the vanishing theorem in cohomology originally obtained by Galicki and Salamon for closed $3$-Sasakian manifolds.
1. Introduction
In [33] Sen, based on predictions coming from a particular type of duality in string theory called $S$-duality, conjectured that the Hodge cohomology $\mathcal{H}^{q}(\widetilde{\mathcal{M}}_{k}^{0})$ of the $L^{2}$-metric of the universal cover ~$\widetilde{\mathcal{M}}_{k}^{0}$ of the reduced moduli space $\mathcal{M}_{k}^{0}$ of $\operatorname{SU}(2)$ monopoles of magnetic charge $k$ on $\mathbb{R}^{3}$ is only non-trivial in middle degree and admits a complete description in terms of a natural $\mathbb{Z}_{k}$-action. Soon after Sen formulated his conjecture, Segal and Selby [31] computed the relative and absolute cohomologies $H_{c}^{q}(\widetilde{\mathcal{M}}_{k}^{0})$ and $H^{q}(\widetilde{\mathcal{M}}_{k}^{0})$ and showed that the images $\operatorname{Im}[H_{c}^{q}(\widetilde{\mathcal{M}}_{k}^{0})\to H^{q}(\widetilde{\mathcal{M}}_{k}^{0})]$ satisfy the predictions of Sen’s conjecture. They used this observation to reformulate Sen’s conjecture into the statement that the natural inclusion
(1.1) $\operatorname{Im}[H_{c}^{q}(\widetilde{\mathcal{M}}_{k}^{0})\to H^{q}(\widetilde{\mathcal{M}}_{k}^{0})]\hookrightarrow\mathcal{H}^{q}(\widetilde{\mathcal{M}}_{k}^{0})$
is in fact an isomorphism.
Around the same time that Sen formulated his conjecture, Nakajima in [29] generalized the $\operatorname{ADHM}$ construction of instantons on asymptotically locally Euclidean $(\operatorname{ALE})$ spaces of [21] by allowing the underying quiver and the dimensions of the vector spaces associated to the vertices of the quiver to be arbitrary. The new family of varieties $\mathfrak{M}_{\zeta}$ thus defined, called Nakajima quiver varieties, carry a natural metric and under the assumption that the $\zeta$ parameter is generic, are complete hyperKähler manifolds.
Shortly after [33, 29] appeared, Vafa and Witten in [35], again based on predictions of $S$-duality, made a similar conjecture to the one of Sen about the Hodge cohomology $\mathcal{H}^{q}(\mathfrak{M}_{\zeta})$ of a quiver variety $\mathfrak{M}_{\zeta}$. Specifically, assuming that the $\zeta$ parameter is generic, they conjectured that the middle dimensional absolute cohomology of $\mathfrak{M}_{\zeta}$ should coincide with all of the Hodge cohomology $\mathcal{H}^{\star}(\mathfrak{M}_{\zeta})$ of the Nakajima metric on $\mathfrak{M}_{\zeta}$. It follows from [30, Corollary 11.2] that the natural map $H_{c}^{q}(\mathfrak{M}_{\zeta})\to H^{q}(\mathfrak{M}_{\zeta})$ from compactly supported into absolute cohomology is an isomorphism in middle degree. Therefore, we can restate the conjecture in the form (1.1) by asking that the natural inclusion
(1.2) $\operatorname{Im}[H_{c}^{q}(\mathfrak{M}_{\zeta})\to H^{q}(\mathfrak{M}_{\zeta})]\hookrightarrow\mathcal{H}^{q}(\mathfrak{M}_{\zeta})$
be in fact an isomorphism.
A major step towards a proof of both conjectures was made by Hitchin [16], who showed that for many hyperKähler metrics coming from hyperKähler quotient constructions, all $L^{2}$-harmonic forms lie in middle dimension, thus immediately proving both conjectures outside of middle degree. The Vafa-Witten conjecture for the simplest type of Nakajima quiver variety, namely the $\operatorname{ALE}$ gravitational instantons, follows from standard results on the $L^{2}$-cohomology of asymptotically conical $(\operatorname{AC})$ metrics, see for example [27, 15]. More recently, attention was focused on the particular case of the Vafa-Witten conjecture when the Nakajima quiver variety is the reduced Hilbert scheme of $n$ points on $\mathbb{C}^{2}$, $\operatorname{Hilb}_{0}^{n}(\mathbb{C}^{2})$. The $n=3$ case of the conjecture was proven by Carron in [8] and in [7] it was proven that the Nakajima metric on $\operatorname{Hilb}_{0}^{n}(\mathbb{C}^{2})$ is quasi-asymptotically locally Euclidean $(\operatorname{QALE})$ in the sense of Joyce [17]. The case for arbitrary $n$ was settled by Kottke and the second author in [18] using the analytical results of [20].
In this paper we study the asymptotic geometry of Nakajima quiver varieties and using the strategy of [18] and the results of [20], we prove that the Vafa-Witten conjecture for Nakajima quiver varieties $\mathfrak{M}_{\zeta}$ with $\zeta$ satisfying a slightly stronger genericity assumption than the one required to guarantee smoothness of $\mathfrak{M}_{\zeta}$.
Theorem 1.1.
The Vafa-Witten conjecture holds for all Nakajima quiver varieties $\mathfrak{M}_{\zeta}$ under the assumption that $\zeta$ is properly generic in the sense of Definition 4.3 below.
In order to apply the results of [20, 18], we need to show that the Nakajima metric on $\mathfrak{M}_{\zeta}$ is a quasi-fibered boundary $(\operatorname{QFB})$ metric in the sense of [9]. In fact, we show that they are quasi-asymptotically conical ($\operatorname{QAC}$), a particular type of $\operatorname{QFB}$ metric of maximal volume growth originally introduced by Degeratu and Mazzeo [11] and generalizing the notion of $\operatorname{QALE}$ metric.
Theorem 1.2.
Given that $\zeta$ is properly generic, the Nakajima metric of any Nakajima quiver variety $\mathfrak{M}_{\zeta}$ is $\operatorname{QAC}$ and admits a smooth expansion at infinity in the sense of Definition 3.4 below. In particular, it is of bounded geometry and of maximal volume growth.
Our strategy to prove this result is strongly inspired by an approach developed by Melrose [25, 23] to give a geometric proof of the result of Carron [7] and show that the metric has a smooth expansion at infinity. We start by radially compactifying the Nakajima quiver representation space $\mathbb{M}$. The action of the group $G$ of gauge transformations on $\mathbb{M}$ is unitary and thus extends to the compactification $\overline{\mathbb{M}}$. Using the result of [2], we resolve the group action at the boundary by iteratively blowing up the boundary strata indexed by conjugacy classes of stabilizer subgroups of $G$. The resulting space $\widetilde{\mathbb{M}}$ is called the $\operatorname{QAC}$ compactification of $\mathbb{M}$. A careful analysis of the hyperKähler moment map $\mu$ shows that the closure $\widetilde{\mu^{-1}(-\zeta)}$ of $\mu^{-1}(-\zeta)$ into $\widetilde{\mathbb{M}}$ is naturally a manifold with fibered corners with induced metric a $\operatorname{QAC}$ metric. Since the whole construction is $G$-equivariant and $G$ acts freely on $\widetilde{\mu^{-1}(-\zeta)}$ the metric descends to a $\operatorname{QAC}$ metric on $\mathfrak{M}_{\zeta}$.
Remark 1.3.
It was proven in [4] that there exist quiver varieties whose associated Nakajima metric is not $\operatorname{QALE}$ and it was asked whether the Nakajima metric on those varieties is $\operatorname{QAC}$. Our result gives a positive answer to their question.
Our analysis of the behaviour of the moment map near the boundary further implies that if $H_{i}$ is a boundary hypersurface of $\widetilde{\mathbb{M}}$ and $\phi_{i}:H_{i}\cap\widetilde{\mu^{-1}(-\zeta)}\to\Sigma_{i}$ is the fiber bundle of the corresponding boundary hypersurface of $\widetilde{\mu^{-1}(-\zeta)}$, then the fiber is the $\operatorname{QAC}$ compactification of a quiver variety of lower dimension. Our assumption that $\zeta$ is properly generic guarantees that the Nakajima quiver varieties appearing as the interiors of these fibers are smooth.
The base $\Sigma_{i}$ turns out to be an incomplete $3$-Sasakian manifold and the induced metric $g_{S_{i}}$ is a $3$-Sasakian exact wedge metric. In order to be able to apply the results of [20], we need to show that the Hodge-deRham operator associated to a flat Euclidean vector bundle $F\to(S_{i},g_{S_{i}})$ is essentially self-adjoint and has some spectral gaps in certain degrees. Specifically, referring to Theorems 5.10 and 5.11 for further details, we prove
Theorem 1.4.
Let $\mathcal{S}$ be a manifold with fibered corners of dimension $34n+$ 3 and $F\to\mathcal{S}$ a flat Euclidean vector bundle. Suppose that $g_{w}$ is an exact wedge $3$-Sasakian metric on $\mathcal{S}$ and let ð$\eth_{w}$ be the Hodge-deRham operator associated to $g_{w}$ and $F$. Then the wedge operator ð$\eth_{w}$ is essentially self-adjoint with self-adjoint extension $vH^{1}_{w}(\mathcal{S};\Lambda^{*}({}^{w}T^{*}\mathcal{S})\otimes F)$ and discrete spectrum. Moreover, for $k\in\{0,\ldots,n\}$,
(1.3) ð$\langle\eth_{w}\psi,\eth_{w}\psi\rangle_{L^{2}_{w}}\geq 2\langle\psi,\psi\rangle_{L^{2}_{w}}\quad\forall\psi\in\Omega^{2k+1}_{c}(\mathcal{S}\setminus\partial\mathcal{S};F).$
In particular, when $\mathcal{S}$ is a closed manifold, this implies and refines the vanishing theorem of Galicki-Salamon [13], namely that the space of harmonic forms in degree $12k+$ 1 is trivial for $k\in\{0,\ldots,n\}$.
The key ingredient behind the proof of this result is the Weitzenböck formula of Semmelmann and Weingart [32] for quaternionic-Kähler metrics.
The paper is organized as follows. In section 2 we review basic facts about Nakajima quiver varieties. In section 3 we introduce manifolds with fibered corners and define $\operatorname{QFB}$ and wedge metrics. In section 4 we construct the $\operatorname{QAC}$ compactification of the Nakajima quiver representation space and use it to show that the natural hyperKähler metric on $\mathfrak{M}_{\zeta}$ is an exact $\operatorname{QAC}$ metric. In section 5 we introduce exact $3$-Sasakian wedge metrics and prove the required mapping properties for the Hodge-deRham operator associated to a flat Euclidean vector bundle $F\to(S,g_{S})$. Finally, in section 6 we use the results of sections 4, 5 along with the results of [20, 18] to prove the Vafa-Witten conjecture.
Acknowledgements.
We are grateful to Vestislav Apostolov and Gilles Carron for pointing out reference [32]. The second author was supported by a NSERC discovery grant and a FRQNT team research project grant.
2. Quiver varieties
Given a finite graph $\Gamma$ with $n$ vertices, let $H$ be the set of pairs consisting of an edge together with an orientation of it. For $h\in H$, let $\operatorname{in}(h)$ denote the incoming vertex, $\operatorname{out}(h)$ the outgoing vertex and $\bar{h}$ the edge with reversed orientation. We will allow for loops, that is, for edges $h$ such that $\operatorname{in}(h)=\operatorname{out}(h)$. Choose a subset $\Omega\subset H$ such that $\overline{\Omega}\cup\Omega=H$ and $\overline{\Omega}\cap\Omega=\emptyset$, where $\overline{\Omega}=\{\overline{h}\;|\;h\in\Omega\}$. Such an $\Omega$ corresponds to a choice of orientation for the graph.
Given a pair of Hermitian vector spaces $(V_{k},W_{k})$ for each vertex $k$, we define the Nakajima quiver representation space
(2.1) $\mathbb{M}(\mathbb{v},\mathbb{w}):=\left(\bigoplus\limits_{h\in H}\operatorname{Hom}(V_{\operatorname{out}(h)},V_{\operatorname{in}(h)})\right)\oplus\left(\bigoplus\limits_{k=1}^{n}\operatorname{Hom}(W_{k},V_{k})\oplus\operatorname{Hom}(V_{k},W_{k})\right),$
where
$\mathbb{v}:=(\dim_{\mathbb{C}}V_{1},...,\dim_{\mathbb{C}}V_{n}),~{}~{}\mathbb{w}:=(\dim_{\mathbb{C}}W_{1},...,\dim_{\mathbb{C}}W_{n}).$
If $B_{h}\in\operatorname{Hom}(V_{\operatorname{out}(h)},V_{\operatorname{in}(h)}),~{}i_{k}\in\operatorname{Hom}(W_{k},V_{k})$ and $j_{k}\in\operatorname{Hom}(V_{k},W_{k})$ then the elements of $\mathbb{M}$ are denoted as triples $(B,i,j)$ where $B$ denotes the collection $(B_{h})_{h\in H}$, $i$ the collection $(i_{k})_{1\leq k\leq n}$ and $j$ the collection $(j_{k})_{1\leq j\leq n}$.
We define on $\mathbb{M}$ the holomorphic symplectic form
(2.2) $\omega_{\mathbb{C}}((B,i,j),(B^{\prime},i^{\prime},j^{\prime})):=\sum\limits_{h\in H}\operatorname{tr}(\epsilon(h)B_{h}B^{\prime}_{\bar{h}})+\sum\limits_{k=1}^{n}\operatorname{tr}(i_{k}j_{k}^{\prime}-i_{k}^{\prime}j_{k}),$
where $\epsilon(h)=1$ if $h\in\Omega$ and $\epsilon(h)=-1$ if $h\in\overline{\Omega}$. The symplectic vector space $\mathbb{M}$ decomposes into the sum $\mathbb{M}=\mathbb{M}_{\Omega}\oplus\mathbb{M}_{\overline{\Omega}}$ of Lagrangian subspaces:
(2.3) $\begin{split}\mathbb{M}_{\Omega}:=&\left(\bigoplus\limits_{h\in\Omega}\operatorname{Hom}(V_{\operatorname{out}(h)},V_{\operatorname{in}(h)})\right)\oplus\left(\bigoplus\limits_{k=1}^{n}\operatorname{Hom}(W_{k},V_{k})\right)\\ \mathbb{M}_{\overline{\Omega}}:=&\left(\bigoplus\limits_{h\in\overline{\Omega}}\operatorname{Hom}(V_{\operatorname{out}(h)},V_{\operatorname{in}(h)})\right)\oplus\left(\bigoplus\limits_{k=1}^{n}\operatorname{Hom}(V_{k},W_{k})\right).\end{split}$
Given this decomposition, we introduce a new complex structure $J$ given by $J\cdot(m,m^{\prime})=(-m^{\prime\dagger},m^{\dagger})$ for $(m,m^{\prime})\in\mathbb{M}_{\Omega}\oplus\mathbb{M}_{\overline{\Omega}}$. The complex structures $I,J$ together endow $\mathbb{M}$ with a hyperKähler structure. Given $\omega_{\mathbb{C}}$ and $J$ we define the metric
(2.4) $g((B,i,j),(B^{\prime},i^{\prime},j^{\prime})):=\omega_{\mathbb{C}}((B,i,j),J\cdot(B^{\prime},i^{\prime},j^{\prime}))=\sum\limits_{h\in H}\operatorname{tr}(B_{h}B_{h}^{{}^{\prime}\dagger})+\sum\limits_{k=1}^{n}\operatorname{tr}(i_{k}i_{k}^{\dagger}+j_{k}^{{}^{\prime}\dagger}j_{k}).$
The group $G=\prod\operatorname{U}(V_{k})$ acts on $\mathbb{M}$. The action, which is given by
$g\cdot(B,i,j)=(g_{\operatorname{in}(h)}B_{h}g_{\operatorname{out}(h)}^{-1},g_{k}i_{k},j_{k}g_{k}^{-1}),$
preserves the metric and the hyperKähler structure. Let $\mu$ be the corresponding hyperKähler moment map. Writing $\mu=(\mu_{\mathbb{R}},\mu_{\mathbb{C}})$ we have that
(2.5) $\begin{split}\mu_{\mathbb{R}}(B,i,j)&=\frac{i}{2}\left(\sum\limits_{h\in H:k=\operatorname{in}(h)}B_{h}B_{h}^{\dagger}-B_{\bar{h}}^{\dagger}B_{\bar{h}}+i_{k}i_{k}^{\dagger}-j_{k}^{\dagger}j_{k}\right)_{k}\in\bigoplus\limits_{k}\mathfrak{u}(V_{k})=\mathfrak{g}\\ \mu_{\mathbb{C}}(B,i,j)&=\left(\sum\limits_{h\in H:k=\operatorname{in}(h)}\epsilon(h)B_{h}B_{\bar{h}}+i_{k}j_{k}\right)_{k}\in\mathfrak{gl}(V_{k})=\mathfrak{g}\otimes\mathbb{C}\end{split}$
Here we identify $\mathfrak{g}$ with its dual $\mathfrak{g}^{\star}$. When the graph $\Gamma$ has loops, we also consider the reduced vector space
$\mathbb{M}_{\operatorname{red}}:=\{(B,i,j)\in\mathbb{M}~{}|~{}\operatorname{Tr}(B_{h})=0\;\mbox{whenever $h$ is a loop}\}\subset\mathbb{M}.$
Notice that the induced metric on $\mathbb{M}_{\operatorname{red}}$ is also hyperKähler and that $\mathbb{M}_{\operatorname{red}}$ is invariant under the action of $G$.
Definition 2.1.
Given an element $\zeta=(\zeta_{\mathbb{R}},\zeta_{\mathbb{C}})\in Z\oplus(Z\otimes\mathbb{C})$, where $Z\subset\mathfrak{g}$ denotes the center, consider the hyperKähler quotient
(2.6) $\mathfrak{M}_{\zeta}:=\{(B,i,j)\in\mathbb{M}~{}|~{}\mu(B,i,j)=-\zeta\}/G.$
This is the quiver variety associated to $H$. If $H$ contains loops, we also consider the reduced quiver variety
(2.7) $\mathfrak{M}_{\zeta}^{\operatorname{red}}:=\{(B,i,j)\in\mathbb{M}_{\operatorname{red}}~{}|~{}\mu(B,i,j)=-\zeta\}/G.$
We want to know when such a quiver variety is smooth. Let $A$ be the adjacency matrix of the graph $\Gamma$, meaning
$a_{i,j}=a_{j,i}=\textrm{number of edges joining vertex }i\textrm{ to vertex }j$
and $C=2\operatorname{Id}-A$ the generalized Cartan matrix. Since the center $Z$ of $\mathfrak{g}$ is the product of scalar matrices on $V_{k}$, we may identify $Z$ with a subspace of $\mathbb{R}^{n}$.
Let
(2.8) $\begin{split}R_{+}:=&\{\theta=(\theta_{k})\in\mathbb{Z}^{n}_{\geq 0}|^{t}\theta C\theta\leq 2\}\backslash\{0\}\\ R_{+}(\mathbb{v}):=&\{\theta\in R_{+}|\theta_{k}\leq\dim_{\mathbb{C}}V_{k}~{}\forall~{}k\}\\ D_{\theta}:=&\left\{x=(x_{k})\in\mathbb{R}^{n}\Big{|}\sum\limits_{k}x_{k}\theta_{k}=0\right\}\textrm{ for }\theta\in R_{+}\\ Z_{s}(\mathfrak{g}):=&\bigcup\limits_{\theta\in R_{+}(\mathbb{v})}\mathbb{R}^{3}\otimes D_{\theta}.\end{split}$
Theorem 2.2 ([29] Theorem 2.8 and Corollary 4.2).
Suppose
(2.9) $\zeta\in\mathbb{R}^{3}\otimes\mathbb{R}^{n}\backslash Z_{s}(\mathfrak{g}).$
Then the quiver variety $\mathfrak{M}_{\zeta}$ is nonsingular and the induced hyperKähler metric is complete. We call parameters satisfying (2.9) generic. Given generic parameters $\zeta\neq\zeta^{\prime}$, $\mathfrak{M}_{\zeta}$ and $\mathfrak{M}_{\zeta^{\prime}}$ are diffeomorphic.
3. Manifolds with fibered corners
Let $M$ be a compact manifold with corners in the sense of [14, 24, 26] and $\mathcal{M}_{1}(M)=\{H_{1},...,H_{l}\}$ be a complete list of its boundary hypersurfaces.
Definition 3.1.
Let $\phi=\{\phi_{1},...,\phi_{l}\}$ be a collection of fiber bundles $\phi_{i}:H_{i}\to S_{i}$ over a compact manifold with corners $S_{i}$. We say that $\phi$ is an iterated fibration structure for $M$ if there exists a partial ordering on the boundary hypersurfaces $H_{i}$ such that:
  • Any subset $\mathcal{I}\subset\{1,...,l\}$ of boundary hypersurfaces such that $\bigcap\limits_{i\in\mathcal{I}}H_{i}\neq\emptyset$ is totally ordered;
  • If $H_{i}<H_{j}$ then $H_{i}\cap H_{j}\neq\emptyset$, $\phi_{i}|_{H_{i}\cap H_{j}}:H_{i}\cap H_{j}\to S_{i}$ is a surjective submersion, $S_{ji}:=\phi_{j}(H_{i}\cap H_{j})$ is a boundary hypersurface of the manifold with corners $S_{j}$ and there is a surjective submersion $\phi_{ji}:S_{ji}\to S_{i}$ such that $\phi_{ji}\circ\phi_{j}=\phi_{i}$ on $H_{i}\cap H_{j}$;
  • The boundary hypersurfaces of $S_{j}$ are given by $S_{ji}$ for $H_{i}<H_{j}$.
In this case, we say that the pair $(M,\phi)$ is a manifold with fibered corners.
For each boundary hypersurface $H_{i}$, let $x_{i}\in C^{\infty}(M)$ be a boundary defining function, that is, $x_{i}$ takes non-negative values, $x_{i}^{-1}(0)=H_{i}$ and $dx_{i}$ is nowhere zero along $H_{i}$. The boundary defining functions $x_{i}$ are compatible with the iterated fibration structure $\phi$ if $x_{i}$ restricted to $H_{j}$ is constant along the fibers of $\phi_{j}:H_{j}\to S_{j}$ whenever $H_{j}>H_{i}$.
Definition 3.2.
Let $v=\prod\limits_{i}x_{i}$ be a total boundary defining function for the manifold with fibered corners $(M,\phi)$. For such a choice, the space $\mathcal{V}_{\operatorname{QFB}}(M)$ of quasi-fibered boundary $(\operatorname{QFB})$ vector fields consists of smooth vector fields $\xi$ in $M$ such that:
  1. (1)
    $\xi$ is tangent to the fibers of $\phi_{i}:H_{i}\to S_{i}$ for each boundary hypersurface $H_{i}$ of $M$;
  2. (2)
    $\xi v\in v^{2}C^{\infty}(M)$.
Remark 3.3.
The condition $\xi v\in v^{2}C^{\infty}(M)$ clearly depends on the choice of the total boundary defining function $v$, but by [20, Lemma 1.1], two total boundary defining functions $v$ and $v^{\prime}$ give the same space of $\operatorname{QFB}$ vector fields if and only if the function $\frac{v}{v^{\prime}}$ is constant on the fibers of $\phi_{i}$ for all $i$. When this is the case, we say that the total boundary defining functions $v$ and $v^{\prime}$ are $\operatorname{QFB}$ equivalent.
As explained in [9], there exists a natural bundle ${}^{\operatorname{QFB}}TM\to M$ called the $\operatorname{QFB}$ tangent bundle and a natural map
(3.1) $a_{\operatorname{QFB}}:^{\operatorname{QFB}}TM\to TM$
inducing a canonical isomorphism
(3.2) $(a_{\operatorname{QFB}})_{\star}:C^{\infty}(M;^{\operatorname{QFB}}TM)\to\mathcal{V}_{\operatorname{QFB}}(M).$
This gives ${}^{\operatorname{QFB}}TM\to M$ a structure of Lie algebroid with anchor map $a_{\operatorname{QFB}}$.
Definition 3.4.
A quasi-fibered boundary $(\operatorname{QFB})$ metric on a manifold with fibered corners $(M,\phi)$ together with a choice of total boundary defining function $v$ is a Riemannian metric on the interior of $M$ of the form
(3.3) $(a_{\operatorname{QFB}})_{\star}(h|_{M\backslash\partial M})$
for some choice of bundle metric $h\in C^{\infty}(M;S^{2}(^{\operatorname{QFB}}T^{\star}M))$ for the vector bundle ${}^{\operatorname{QFB}}TM$. In this case we say that the manifold with corners $M$ is the $\operatorname{QFB}$ compactification of the corresponding Riemannian manifold.
A $\operatorname{QFB}$ metric on $(M,\phi)$ such that for each maximal hypersurface $H_{i}$, $S_{i}=H_{i}$ and $\phi_{i}=\operatorname{Id}$ will be called a quasi-asymptotically conical $(\operatorname{QAC})$ metric. Two $\operatorname{QFB}$ respectively $\operatorname{QAC}$ metrics on $M$ $g,g^{\prime}$ are called quasi-isometric if there exists a positive constant $C$ such that
$\frac{g}{C}<g^{\prime}<Cg$
everywhere on $M$.
For the purposes of the paper, we introduce two more types of metrics on manifolds with fibered corners. If we drop condition $(2)$ from definition 3.2 we obtain the Lie algebra of edge vector fields $\mathcal{V}_{e}(M)$ and the edge tangent bundle ${}^{e}TM\to M$, a naturally associated Lie algebroid with anchor map $a_{e}:{}^{e}TM\to TM$. As with $\mathcal{V}_{\operatorname{QFB}}(M)$, to the Lie algebra $\mathcal{V}_{e}(M)$ one can associate the class of edge metrics on $M\backslash\partial M$. Both $\operatorname{QFB}$ metrics $(M\backslash\partial M,g_{\operatorname{QFB}})$ and edge metrics $(M\backslash\partial M,g_{e})$ are examples of Riemannian manifolds with Lie structure at infinity in the sense of [3]. By [3] and [6], such metrics are complete metrics of infinite volume with bounded geometry. Moreover, by [3], all metrics corresponding to a fixed Lie structure at infinity are quasi-isometrics.
In order to introduce wedge metrics, we define the wedge cotangent bundle
(3.4) ${}^{w}T^{\star}M:=v(^{e}T^{\star}M)$
with $v$ a total boundary defining function.
Definition 3.5.
A wedge metric (also called incomplete iterated edge metrics in [1]) on a manifold with fibered corners $(M,\phi)$ is a Riemannian metric $g_{w}$ on the interior of $M$ of the form
(3.5) $g_{w}=a_{\star}(h|_{M\backslash\partial M})$
for some choice of bundle metric $h\in C^{\infty}(M;S^{2}(^{w}T^{\star}M))$. Equivalently, a wedge metric is a metric of the form $g_{w}=v^{2}g_{e}$ for some edge metric $g_{e}$.
A wedge metric $g_{w}$ is of finite volume and is geodesically incomplete, so the pair $(M\backslash\partial M,g_{w})$ is not a Riemannian manifold with Lie structure at infinity. This is evident from the fact that the wedge vector fields $\xi\in{{}^{w}TM}$ are not preserved by the Lie bracket.
In this paper, we will need to use exact wedge metrics in the sense of [20, Definition 8.4]. To explain what this is, we need first to recall what are wedge metrics of product type. Let $c_{i}:H_{i}\times[0,\delta_{i})\to M$ be a collar neighbourhood of $H_{i}$ compatible with the boundary defining functions in the sense that $c_{i}^{\star}x_{i}$ corresponds to the projection $\operatorname{pr}_{2}:H_{i}\times[0,\delta_{i})\to[0,\delta_{i})$ and $c_{i}^{\star}x_{j}$ is the pullback of a function on $H_{i}$. Choose a connection on the fiber bundle $\phi_{i}:H_{i}\to S_{i}$. On $S_{i}\backslash\partial S_{i}$ let $g_{S_{i}}$ be a wedge metric compatible with the iterated fibration structure on $S_{i}$. Let $\kappa_{i}$ be a family of fiberwise wedge metrics in the fibers of $\phi_{i}:H_{i}\to S_{i}$. Using the connection on $\phi_{i}$, this family can be lifted to a vertical symmetric $2$-tensor on $H_{i}\backslash\partial H_{i}$. Let $\operatorname{pr}_{1}:H_{i}\times[0,\delta_{i})\to H_{i}$ be the projection onto the first factor. A product type wedge metric near $H_{i}$ is given by
(3.6) $g_{w}=\rho_{i}^{2}(dx_{i}^{2}+\operatorname{pr}_{1}^{\star}\phi_{i}^{\star}g_{S_{i}}+x_{i}^{2}\operatorname{pr}_{1}^{\star}\kappa_{w,i}),\quad with\quad\rho_{i}=\prod_{H_{j}<H_{i}}x_{j},$
where $g_{S_{i}}$ is an edge metric on $S_{i}$ and $\kappa_{w,i}=\left(\frac{v_{i}^{2}}{x_{i}^{2}}\right)\kappa_{i}$ is a $2$-tensor inducing a wedge metric on the fibers of $\phi_{i}:H_{I}\to S_{i}$ in such a way that $\phi_{i}^{*}g_{S_{i}}+\kappa_{w,i}$ is a Riemmannian metric turning $\phi_{i}$ into a Riemannian submersion onto $(S_{i},g_{S_{i}})$. Notice in particular that $g_{S_{i},w}:=\rho_{i}^{2}g_{S_{i}}$ is the natural wedge metric induced by $g_{w}$ on $S_{i}$.
Definition 3.6.
An exact wedge metric is a wedge metric which is of product type near $H_{i}$ up to a term in $x_{i}C^{\infty}(M;S^{2}(^{w}T^{\star}M))$ for each boundary hypersurface $H_{i}$ of $M$.
There is also a notion of product type and exact $\operatorname{QFB}$ metrics. The definition of a product type $\operatorname{QFB}$ metric is very similar to the one of a product wedge metric, however this time instead of the level sets of $x_{i}$ we use the level sets of some total boundary defining function $\prod_{j}x_{j}$ $\operatorname{QFB}$ equivalent to the total boundary defining function $v$ defining $\mathcal{V}_{\operatorname{QFB}}(M)$. To indicate that we could take different total boundary defining functions compatible with the Lie algebra of $\operatorname{QFB}$ vector fields fo different boundary hypersurfaces, we will denote this total boundary defining function by $u_{i}$. With this understood, we consider the open set
(3.7) $\mathcal{U}_{i}=\{(p,\tau)\in H_{i}\times[0,\delta_{i})|\prod\limits_{j\neq i}x_{j}(p)>\frac{\tau}{\delta_{i}}\}\subset H_{i}\times[0,\delta_{i})$
with natural diffeomorphism
(3.8) $\begin{split}\psi_{i}:(H_{i}\backslash\partial H_{i})&\times[0,\delta_{i})\to\mathcal{U}_{i}\\ (&p,t)\longrightarrow(p,t\prod\limits_{j\neq i}x_{j}(p)).\end{split}$
Following the definition of a product wedge metric, we choose a connection for the fiber bundle $\phi_{i}:H_{i}\to S_{i}$ and let $g_{S_{i}}$ be a wedge metric. Let $\kappa_{i}$ be a family of fiberwise $\operatorname{QFB}$ metrics in the fibers of $\phi_{i}:H_{i}\to S_{i}$ and use the connection to lift them to a vertical symmetric $2$-tensor on $H_{i}\backslash\partial H_{i}$. In $\mathcal{U}_{i}$ seen as a subset of $H_{i}\times[0,\delta_{i})$, a $\operatorname{QFB}$ metric of product type near $H_{i}$ is a metric of the form
(3.9) $g_{\operatorname{QFB}}=\frac{\,du_{i}^{2}}{u_{i}^{4}}+\frac{\operatorname{pr}_{1}^{\star}\phi_{i}^{\star}g_{S_{i}}}{u_{i}^{2}}+\operatorname{pr}_{1}^{\star}\kappa_{i}.$
More generally, an exact $\operatorname{QFB}$ metric is a $\operatorname{QFB}$ metric which is of product type near each hypersurface $H_{i}$ up to a term in $x_{i}C^{\infty}(M;S^{2}(^{\operatorname{QFB}}T^{\star}M))$.
4. Compactification of Nakajima quiver representation spaces
For $\zeta=0$, the quiver variety $\mathfrak{M}_{\zeta}$ is singular. However, as described in [29], it has a natural structure of a stratified space. Let $\mathbb{q}:\mathbb{M}\to\mathbb{M}/G$ be the natural quotient map.
Lemma 4.1.
A point $m\in\mu^{-1}(-\zeta)$ is a singular point of $\mu^{-1}(-\zeta)$ if and only if $\mathbb{q}(m)\in\mathfrak{M}_{\zeta}$ is a singular point of $\mathfrak{M}_{\zeta}$.
Proof.
If $m\in\mu^{-1}(-\zeta)$ is singular, then the differential of $\mu$ at $m$ is not surjective. By the definition of a hyperKähler moment map, this means that the stabilizer of $m$ is non-trivial, so the corresponding point $x\in\mu^{-1}(-\zeta)/G$ is singular. Conversely, if $\mathbb{q}(m)\in\mu^{-1}(-\zeta)/G$ is singular, then the stabilizer $G_{m}$ of $m$ is non-trivial. By [29, p.391], this can only happen if the Lie algebra $\mathfrak{g}_{m}$ of $G_{m}$ is non-trivial. By the definition of a hyperKähler moment map, this means that $d\mu$ is not surjective at $m$, hence that $\mu^{-1}(-\zeta)$ is not smooth at $m$. ∎
More generally, the quotient $\mathbb{M}/G$ is singular, but it is a smoothly stratified space by [2] with strata given by
$\mathbb{M}^{I}/G=\{m\in\mathbb{M}\;|\;G_{m}\in I\}/G$
for $I$ a conjugacy class of a stabilizer subgroup. Strictly speaking, the result of [2] is for compact manifolds with corners, but since the action of $G$ naturally extends to the radial compactification $\overline{\mathbb{M}}$ of $\mathbb{M}$, it suffices to apply the result of [2] to $\overline{\mathbb{M}}$ and restrict it to $\mathbb{M}$. Even if $\mathbb{M}$ itself is smooth, this induces a corresponding structure of smoothly stratified space with strata given by
$\mathbb{M}^{I}=\{m\in\mathbb{M}\;|\;G_{m}\in I\}$
for $I$ a conjugacy class of a stabilizer subgroup. The regular stratum is the one corresponding to the conjugacy class of the trivial stabilizer $G_{m}=\{\operatorname{Id}\}$, while the deepest stratum is the origin and corresponds to the conjugacy class of the stabilizer $G_{m}=G$. Thus, if $m\in\mathbb{M}\setminus\{0\}$ is such that $\mathbb{q}(m)\in\mathbb{M}/G$ is singular, its stabilizer $G_{m}$ is non-trivial and is strictly contained in $G$. Let $T_{m}\mathcal{O}_{m}\subset T_{m}\mathbb{M}$ be the tangent space at $m$ of the orbit $\mathcal{O}_{m}$ of $m$. Let $\widehat{\mathbb{M}}_{m}$ be the orthogonal complement in $T_{m}\mathbb{M}$ to the $\mathbb{H}$-module
$\mathbb{H}T_{m}\mathcal{O}_{m}=T_{m}\mathcal{O}_{m}\oplus I_{1}T_{m}\mathcal{O}_{m}\oplus I_{2}T_{m}\mathcal{O}_{m}\oplus I_{3}T_{m}\mathcal{O}_{m}.$
Clearly, $\widehat{\mathbb{M}}_{m}$ is itself an $\mathbb{H}$-module and the action of $G$ on $\mathbb{M}$ induces an action of $G_{m}$ on $\widehat{\mathbb{M}}_{m}$. Using the canonical identification of $T_{m}\mathbb{M}$ with $\mathbb{M}$, $\widehat{\mathbb{M}}_{m}$ can be seen as an $\mathbb{H}$-submodule of $\mathbb{M}$. Moreover, since the action of $G$ commutes with the natural $\mathbb{R}^{+}$ action on $\mathbb{M}$, there is a canonical identification $\widehat{\mathbb{M}}_{\lambda m}=\widehat{\mathbb{M}}_{m}$ for any $\lambda\in\mathbb{R}^{+}$.
Lemma 4.2.
Fix $m\in\mu^{-1}(0)\backslash\{0\}\subset\mathbb{M}\backslash\{0\}$ with non-trivial stabilizer $G_{m}$. The restriction $\hat{\mu}_{m}$ of the moment map $\mu$ to $\widehat{\mathbb{M}}_{m}$ induces a hyperKähler moment map with values in $\mathfrak{g}^{*}_{m}$ for the action of $G_{m}$ on $\widehat{\mathbb{M}}_{m}$, where $\mathfrak{g}_{m}$ is the Lie algebra of $G_{m}$.
Proof.
Indeed, since
(4.1) $\langle d\mu_{i}(v),\xi\rangle=\omega_{i}(\xi_{*},v)=g(I_{i}\xi_{*},v)\quad\forall\xi\in\mathfrak{g},$
where $\xi_{*}\in\mathcal{C}^{\infty}(\mathbb{M};T\mathbb{M})$ is the vector field generated by $\xi$, we see that
(4.2) $\langle d\mu,\xi\rangle|_{\widehat{\mathbb{M}}_{m}}=0$
for $\xi\in(\mathfrak{g}_{m}^{\perp})^{*}$, where $\mathfrak{g}^{\perp}_{m}$ is the orthogonal complement of $\mathfrak{g}_{m}$ in $\mathfrak{g}$. Since $m\in\mu^{-1}(0)$, notice that $m\in\widehat{\mathbb{M}}_{m}$ where $\widehat{\mathbb{M}}_{m}$ is seen as a subspace of $\mathbb{M}$. This observation, together with the vanishing of the the derivative (4.2) implies that
$\langle\mu,\xi\rangle|_{\widehat{\mathbb{M}}_{m}}=0\quad\forall\xi\in(\mathfrak{g}_{m}^{\perp})^{*}.$
Hence, the restriction of $\mu$ to $\widehat{\mathbb{M}}_{m}$ takes values in $\mathfrak{g}_{m}^{\star}$ and therefore corresponds to the hyperKähler moment map of the action of $G_{m}$ on $\widehat{\mathbb{M}}_{m}$. ∎
Now, recall from [29, Lemma 6.5] that since $m\in\mu^{-1}(0)\backslash\{0\}\subset\mathbb{M}\backslash\{0\}$ has non-trivial stabilizer $G_{m}$, there is an induced orthogonal decomposition
(4.3) $V=V^{(0)}\oplus(V^{(1)})^{\oplus\hat{v}_{1}}\oplus\cdots\oplus(V^{(r)})^{\oplus\hat{v}_{r}}$
of $V=\oplus_{k=1}^{n}V_{k}$. On $\widehat{\mathbb{M}}_{m}$, this induces a decomposition
(4.4) $\widehat{\mathbb{M}}_{m}=(\widehat{\mathbb{M}}_{m}\cap\mathbb{M}(v^{(0)},w))\oplus\left({\bigoplus_{i,j=1}^{r}\widehat{M}_{ij}\otimes\operatorname{Hom}(\mathbb{C}^{\hat{v}_{i}},\mathbb{C}^{\hat{v}_{j}})}\right)\oplus\left({\bigoplus_{i=1}^{r}\operatorname{Hom}(\mathbb{C}^{\hat{v}_{i}},\widehat{W}_{i})}\right)\oplus\left({\bigoplus_{i=1}^{r}\operatorname{Hom}(\widehat{W}_{i},\mathbb{C}^{\hat{v}_{i}})}\right),$
where
$\widehat{M}_{ij}:=\widehat{\mathbb{M}}\cap\left({\bigoplus_{h\in H}\operatorname{Hom}(V^{(i)}_{\operatorname{out}(h)},V^{(i)}_{\operatorname{in}(h)})}\right)$
and
$\widehat{W}_{i}:=\widehat{\mathbb{M}}\cap\left\{\left({\bigoplus_{h\in H}\operatorname{Hom}(V^{(0)}_{\operatorname{out}(h)},V^{(i)}_{\operatorname{in}(h)})}\right)\oplus\left({\bigoplus_{k=1}^{n}\operatorname{Hom}(W_{k},V^{(i)}_{k})}\right)\right\}.$
In this decomposition, the stabilizer $G_{m}$ is given by
(4.5) $G_{m}=\prod_{i=1}^{r}\operatorname{U}(\hat{v}_{i})$
and acts trivially on the component
(4.6) $\mathbb{T}_{m}:=(\widehat{\mathbb{M}}_{m}\cap\mathbb{M}(v^{(0)},w))\oplus\left({\bigoplus_{i=1}^{r}\widehat{M}_{ii}\otimes\operatorname{Id}_{\mathbb{C}^{\hat{v}_{i}}}}\right).$
Notice that $(\oplus_{i=1}^{3}I_{i}T_{m}\mathcal{O}_{m})\oplus\mathbb{T}_{m}$ can be identified with the tangent space at $\mathbb{q}(m)$ of the stratum of $M/G$ containing $\mathbb{q}(m)$. Let $\mathbb{T}_{m}^{\perp}$ be the orthogonal complement of $\mathbb{T}_{m}$ in $\widehat{\mathbb{M}}_{m}$. Since the action of $G_{m}$ is trivial on $\mathbb{T}_{m}$, the moment map $\hat{\mu}_{m}$ is trivial on $\mathbb{T}_{m}$, that is, it factors through the projection
$\widehat{\mathbb{M}}_{m}=\mathbb{T}_{m}\oplus\mathbb{T}_{m}^{\perp}\to\mathbb{T}_{m}^{\perp}$
and can be seen as a moment map on $\mathbb{T}_{m}^{\perp}$,
(4.7) $\hat{\mu}_{m}:\mathbb{T}^{\perp}_{m}\to\mathbb{R}^{3}\otimes\mathfrak{g}_{m}^{*}.$
In fact, $\mathbb{T}^{\perp}_{m}$ is naturally a reduced Nakajima quiver representation space, namely
(4.8) $\mathbb{T}_{m}^{\perp}=\mathbb{M}(\hat{v},\hat{w})_{\operatorname{red}}$
with $\hat{v}={}^{t}(\hat{v}_{1},\ldots,\hat{v}_{r})$, $\hat{w}={}^{t}(\dim_{\mathbb{C}}\widehat{W}_{1},\ldots,\dim_{\mathbb{C}}\widehat{W}_{r})$ and with adjacency matrix $\hat{A}$ of the graph having $(i,j)$-component $\dim_{\mathbb{C}}\widehat{M}_{ij}$. From the description above, $\mathbb{T}^{\perp}_{m}$ is the orthogonal complement of the tangent space at $m$ of the stratum of $\mathbb{M}$ containing $m$. Clearly, the natural $\mathbb{R}^{+}$-action on $\mathbb{M}$ induces a canonical identification
$\mathbb{T}^{\perp}_{\lambda m}=\mathbb{T}^{\perp}_{m}\quad\forall\lambda\in\mathbb{R}^{+},\;\forall m\in\mathbb{M}\setminus\{0\}.$
For $\zeta\in\mathbb{R}^{3}\otimes Z\subset\mathbb{R}^{3}\otimes\mathfrak{g}^{*}$, denote by $\zeta_{m}$ its image under the canonical map
$\operatorname{pr}_{m}:\mathbb{R}^{3}\otimes\mathfrak{g}^{*}\to\mathbb{R}^{3}\otimes\mathfrak{g}^{*}_{m}.$
Obvioulsy, $\zeta_{m}$ is in the center of $\mathfrak{g}_{m}^{*}$, so $G_{m}$ acts on the preimage $\hat{\mu}_{m}^{-1}(-\zeta_{m})$ of the moment map (4.7). By the discussion above, the quotient $\hat{\mu}^{-1}(-\zeta_{m})/G_{m}$ is a (reduced) Nakajima quiver variety.
In order for the arguments of sections 5 and 6 to work, we need to assume that $\hat{\mu}^{-1}(-\zeta_{m})/G_{m}$ is smooth for any choice of $m\in\mu^{-1}(0)\backslash\{0\}\subset\mathbb{M}\backslash\{0\}$ with non-trivial stabilizer $G_{m}$. From Theorem 2.2 this condition is equivalent to requiring that $\zeta\notin Z_{s}(\mathfrak{g})$ and that $\forall m\in\mu^{-1}(0)\backslash\{0\}$ such that $G_{m}\neq\{\operatorname{Id}\}$, $\zeta_{m}\notin Z_{s}(\mathfrak{g}_{m})$.
Definition 4.3.
We call $\zeta$ properly generic if it satisfies the above condition.
Lemma 4.4.
The subset of $\mathbb{R}^{3}\otimes Z$ where $\zeta$ fails to be properly generic is of real co-dimension three.
Proof.
We first need to properly define the sets $Z_{s}(\mathfrak{g}_{m})$ as we did for the center of the full Lie algebra $\mathfrak{g}$ in (2.8). Given $m$ as above, the Cartan matrix is given by $C_{m}=2\operatorname{Id}-\hat{A}_{m}$ with $\hat{A}_{m}$ the adjacency matrix of the reduced quiver variety (4.8), while $R_{+}(\mathbb{v})$ is replaced by $R_{+}(\hat{v})$. By (4.5) and [29, Lemma 6.5 (6)] , the center of $G_{m}$ is $Z_{m}=\prod_{i=1}^{r}U_{m}^{i}(1)$ with
$U_{m}^{i}(1):=G_{m}\cap\prod\limits_{k}U(V_{k}^{(i)})\cong U(1).$
Its Lie algebra is therefore given by $\mathfrak{z}_{m}=\bigoplus_{i=1}^{r}\mathfrak{u}_{m}^{i}(1)$ with $\mathfrak{u}_{m}^{i}(1)$ the Lie algebra of $U_{m}^{i}(1)$. Consequently, $Z_{s}(\mathfrak{g}_{m}):=\bigcup_{\theta\in R_{+}(\hat{v})}\mathbb{R}^{3}\otimes D^{m}_{\theta}$ with
$D^{m}_{\theta}:=\left\{x=(x_{i})\in\bigoplus_{i=1}^{r}\mathfrak{u}^{i}_{m}(1)\Big{|}\sum\limits_{k}x_{k}\theta_{k}=0\right\}\textrm{ for }\theta\in R_{+}(\hat{v}).$
As is clear from (4.5) and the construction of the decomposition (4.3), the subset $Z_{s}(\mathfrak{g}_{m})$ only depends on $m$ through its stabilizer $G_{m}$, namely $Z_{s}(\mathfrak{g}_{m})=Z_{s}(\mathfrak{g}_{m^{\prime}})$ whenever $\mathfrak{g}_{m}=\mathfrak{g}_{m^{\prime}}$.
However, in a given conjugacy class of stabilizer subgroups, the subset $Z_{s}(\mathfrak{g}_{m})$ may vary, in particular along the orbit $\mathcal{O}_{m}$ of $m$. Nevertheless, along such an orbit, the conditions $\zeta_{m^{\prime}}\in Z_{s}(\mathfrak{g}_{m^{\prime}})$ for $m^{\prime}\in\mathcal{O}_{m}$ only correspond to one condition, since by construction,
(4.9) $\zeta_{m}\in Z_{s}(\mathfrak{g}_{m})\Leftrightarrow\zeta_{gm}\in Z_{s}(\mathfrak{g}_{gm})\quad\forall g\in G.$
Indeed, since moving from $m$ to $gm$ changes the whole decomposition (4.3) by its composition by $g$, the Cartan matrix and $R_{+}(\hat{v})$ are the same for $m$ and $gm$. Since $G_{gm}=gG_{m}g^{-1}$, it follows that $\mathfrak{u}_{gm}^{i}(1)=g\mathfrak{u}_{m}^{i}(1)g^{-1}$ for all $i$ and therefore that $D^{gm}_{\theta}=gD^{m}_{\theta}g^{-1}$ and $Z_{s}(\mathfrak{g}_{gm})=gZ_{s}(\mathfrak{g}_{m})g^{-1}$. Hence, (4.9) follows from the fact that $\mathfrak{g}_{gm}=g\mathfrak{g}_{m}g^{-1}$, so in particular $\zeta_{gm}=g\zeta_{m}g^{-1}$. As a result, the condition that $\zeta_{m}\in Z_{s}(\mathfrak{g}_{m})$ only depends on the conjugacy class of $G_{m}$.
Since there are only finitely many conjugacy classes of stabilizer subgroups in $G$, the result follows provided we can show that $\operatorname{pr}_{m}^{-1}(\mathbb{R}^{3}\otimes Z_{s}(\mathfrak{g}_{m}))$ is of real co-dimension three inside $\mathbb{R}^{3}\otimes Z$. To see this, let $\mathfrak{u}(1)$ and $\mathfrak{u}_{m}(1)$ denote the Lie algebras of scalars for $G$ and $G_{m}$. By [29, Lemma 6.5 (6)], the restriction of $\operatorname{pr}_{m}$ to $\mathbb{R}^{3}\otimes\mathfrak{u}(1)\subset\mathbb{R}^{3}\otimes Z$ induces a surjective map
$\operatorname{pr}_{m}:\mathbb{R}^{3}\otimes\mathfrak{u}(1)\to\mathbb{R}^{3}\otimes\mathfrak{u}_{m}(1).$
On the other hand, by (2.8), an element $\zeta_{m}\in\mathbb{R}^{3}\otimes\mathfrak{u}_{m}(1)$ is in $Z_{s}(\mathfrak{g}_{m})$ if and only if $\zeta_{m}=0$. This shows that $\mathbb{R}^{3}\otimes Z_{s}(\mathfrak{g}_{m})\cap\operatorname{pr}_{m}(\mathbb{R}^{3}\otimes Z)$ is of real codimension three in $\operatorname{pr}_{m}(\mathbb{R}^{3}\otimes Z)$, hence that $\operatorname{pr}_{m}^{-1}(\mathbb{R}^{3}\otimes Z_{s}(\mathfrak{g}_{m}))$ is of real co-dimension three inside $\mathbb{R}^{3}\otimes Z$ has desired.
We can now introduce the natural compactification of $\mathbb{M}$ to describe the asymptotic geometry at infinity of the associated quiver variety. Let us first denote by $\overline{\mathbb{M}}$ the radial compactification of $\mathbb{M}$. Let $\mathfrak{s}_{1},\ldots,\mathfrak{s}_{\ell}$ the strata of $\mathbb{S}(\mathbb{M})=\partial\overline{\mathbb{M}}$ as a $G$-manifold listed in an order compatible with the partial order, namely
$\mathfrak{s}_{i}<\mathfrak{s}_{j}\;\Longrightarrow\;i<j.$
In particular, $\mathfrak{s}_{\ell}$ corresponds to the regular stratum of $\partial\overline{\mathbb{M}}$.
Definition 4.5.
The $\operatorname{QAC}$ compactification of $\mathbb{M}$ is the manifold with corners
$\widetilde{\mathbb{M}}:=[\overline{\mathbb{M}};\overline{\mathfrak{s}}_{1},\ldots,\overline{\mathfrak{s}}_{\ell-1}].$
In this defintion, the order in which we blow-up is important. First, since $\mathfrak{s}_{1}$ is minimal with respect to the partial order, $\mathfrak{s}_{1}=\overline{\mathfrak{s}}_{1}$ is a closed submanifold of $\partial\overline{M}$, so its blow-up is well-defined, as well as the blow-ups of all minimal strata. More generally, before the blow-up of $\overline{s}_{i}$ is performed, notice that $\overline{s}_{j}$ has been blown up whenever $\mathfrak{s}_{j}<\mathfrak{s}_{i}$, so by [2, Proposition 7.4 and Theorem 7.5], the lift of $\overline{s}_{i}$ is a $p$-submanifold and its blow-up is well-defined. The manifold with corners $\widetilde{\mathbb{M}}$ has $\ell$ boundary hypersurfaces $H_{1},\ldots,H_{\ell}$ corresponding to the strata $\mathfrak{s}_{1},\ldots,\mathfrak{s}_{\ell}$ of $\partial\overline{\mathbb{M}}$. By [2, Theorem 7.5], the maximal hypersurface $H_{\ell}$ has an iterated fibration structure, namely it is a manifold with fibered corners. Clearly, this iterated fibration naturally extends to induce on $\widetilde{M}$ an iterated fibration structure with fiber bundle
(4.10) $\phi_{i}:H_{i}\to S_{i}$
induced by the blow-down map
$\widetilde{\mathbb{M}}\to[\overline{\mathbb{M}};\overline{\mathfrak{s}}_{1},\ldots,\overline{\mathfrak{s}}_{i-1}]$
with $S_{i}$ the manifold with fibered corners resolving the smoothly stratified space $\overline{s}_{i}$. To describe the fibers of (4.10), consider first the case $i=1$ and let $\overline{H}_{1}$ be the boundary hypersurface in $[\overline{\mathbb{M}};\overline{\mathfrak{s}_{1}}]$ created by the blow-up of $\overline{s}_{1}$. Then the blow-down map $[\overline{\mathbb{M}};\overline{\mathfrak{s}}_{1}]\to\overline{\mathbb{M}}$ induces a fiber bundle
$\overline{\phi}_{1}:\overline{H_{1}}\to S_{1}=\overline{s}_{1}.$
By the discussion above, the fibers above $m\in S_{1}$ is the radial compactification $\overline{\mathbb{T}^{\perp}_{m}}$ of the space orthogonal to the tangent space of the stratum of $\partial\overline{\mathbb{M}}$ containing $m$. Now, when we perform the blow-ups of the other strata of $\partial\overline{\mathbb{M}}$, this corresponds on $\overline{\mathbb{T}^{\perp}_{m}}$ to the blow-up of the intersection of lifts of the strata $\overline{\mathfrak{s}}_{2},\ldots\overline{\mathfrak{s}}_{\ell-1}$ with $\partial\overline{\mathbb{T}^{\perp}_{m}}$, so that on $\widetilde{\mathbb{M}}$, $\overline{\mathbb{T}^{\perp}_{m}}$ lifts to the its $\operatorname{QAC}$ compactification $\widetilde{\mathbb{T}^{\perp}_{m}}$. For the other boundary hypersurfaces of $\widetilde{\mathbb{M}}$, the same phenomenon occurs if $m\in S_{i}\setminus\partial S_{i}$, namely the fiber above $m$ corresponds to the $\operatorname{QAC}$ compactification $\widetilde{\mathbb{T}^{\perp}_{m}}$ of $\mathbb{T}^{\perp}_{m}$. When $H_{i}$ is not minimal and $m\in\partial S_{i}$, this is also what happens since $\mathbb{T}^{\perp}_{m}$ is now the orthogonal of complement of $T_{m}S_{i}$ in the tangent space at $m$ in
$[\partial\overline{\mathbb{M}};\overline{\mathfrak{s}}_{1},\ldots,\overline{\mathfrak{s}}_{i-1}]$
with $G_{m}$ the stabilizer of $m$ in $[\partial\overline{\mathbb{M}};\overline{\mathfrak{s}}_{1},\ldots,\overline{\mathfrak{s}}_{i-1}]$.
For $\zeta\in\mathbb{R}^{3}\otimes Z$, let $\overline{\mu^{-1}(-\zeta)}$ and $\widetilde{\mu^{-1}(-\zeta)}$ be the closure of $\mu^{-1}(\zeta)$ inside $\overline{\mathbb{M}}$ and $\widetilde{\mathbb{M}}$ respectively. Since $\mu$ is homogeneous of degree $2$ with respect to the natural $\mathbb{R}^{+}$-action on $\mathbb{M}$, notice that $\mu^{-1}(0)$ is a cone in $\mathbb{M}$ with possibly a singular cross-section, while for $\zeta\in\mathbb{R}^{3}\otimes Z$ fixed, $\mu^{-1}(-\zeta)$ is asymptotic to $\mu^{-1}(0)$ at infinity in the sense that
(4.11) $\overline{\mu^{-1}(-\zeta)}\cap\partial\overline{\mathbb{M}}=\overline{\mu^{-1}(0)}\cap\partial\overline{\mathbb{M}}.$
Hence, unless the cone $\mu^{-1}(0)$ has a smooth cross-section, even if $\mu^{-1}(-\zeta)$ is smooth, for instance when $\zeta$ is generic, $\overline{\mu^{-1}(-\zeta)}$ will not be smooth and will have singularities on its boundary $\partial\overline{\mu^{-1}(-\zeta)}=\overline{\mu^{-1}(0)}\cap\partial\overline{\mathbb{M}}$. Now, by Lemma 4.1, the cone $\mu^{-1}(0)$ is naturally a stratified space with strata induced by those of the $G$-space $\mathbb{M}$. Similarly, $\overline{\mu^{-1}(0)}$ and $\partial\overline{\mu^{-1}(0)}=\overline{\mu^{-1}(0)}\cap\partial\overline{\mathbb{M}}$ are smoothly stratified spaces with strata induced by those of $\overline{\mathbb{M}}$ and $\partial\overline{\mathbb{M}}$. In particular, for each stratum $\mathfrak{s}_{i}\subset\partial\overline{\mathbb{M}}$, there is a corresponding stratum $\mathfrak{s}_{i}\cap\partial\overline{\mu^{-1}(0)}$ of $\partial\overline{\mu^{-1}(0)}$. On $\widetilde{\mathbb{M}}$, we can correspondingly associate to each $H_{i}\in\mathcal{M}_{1}(\widetilde{\mathbb{M}})$ a ‘boundary hypersurface’ $H_{i}\cap\widetilde{\mu^{-1}(0)}$ of $\widetilde{\mu^{-1}(0)}$ with
(4.12) $\Sigma_{i}:=\phi_{i}(H_{i}\cap\widetilde{\mu^{-1}(0)})\subset S_{i}$
a manifold with with fibered corners corresponding to a resolution of the closed stratum $\overline{\mathfrak{s}}_{i}\cap\partial\overline{\mu^{-1}(0)}$.
Theorem 4.6.
For $\zeta\in\mathbb{R}^{3}\otimes Z$ properly generic, $\widetilde{\mu^{-1}(-\zeta)}$ is naturally a manifold with fibered corners. For each $H_{i}\in\mathcal{M}_{1}(\widetilde{\mathbb{M}})$ it has a boundary hypersurface $H_{i}\cap\widetilde{\mu^{-1}(-\zeta)}$ with fiber bundle
(4.13) $\phi_{i}:H_{i}\cap\widetilde{\mu^{-1}(-\zeta)}\to\Sigma_{i}$
induced by restriction of the fiber bundle $\phi_{i}:H_{i}\to S_{i}$. Moreover, the free $G$-action on $\mu^{-1}(-\zeta)$ extends to a free $G$-action on $\widetilde{\mu^{-1}(-\zeta)}$ as a manifold with fibered corners.
Proof.
Proceeding by induction on the depth of the Nakajima quiver representation space $\mathbb{M}$ as a $G$-space, we can assume that the result hold for Nakajima quiver representation spaces of lower depth. Let us first consider the case where $H_{i}$ is minimal. Given $m\in\Sigma_{i}\subset S_{i}$, the corresponding fiber $\phi_{i}^{-1}(m)$ in $H_{i}$ is the $\operatorname{QAC}$ compactification $\widetilde{\mathbb{T}^{\perp}_{m}}$ of a certain Nakajima quiver representation space of lower depth. In terms of the decomposition
(4.14) $\mathbb{R}^{3}\otimes\mathfrak{g}^{*}=(\mathbb{R}^{3}\otimes\mathfrak{g}_{m}^{*})\oplus(\mathbb{R}^{3}\otimes(\mathfrak{g}_{m}^{\perp})^{*})$
the moment $\mu$ has a decomposition
(4.15) $\mu=\hat{\mu}_{m}+\check{\mu}_{m}$
with
(4.16) $\hat{\mu}_{m}:\mathbb{M}\to\mathbb{R}^{3}\otimes\mathfrak{g}_{m}^{*}\quad\mbox{and}\quad\check{\mu}_{m}:\mathbb{M}\to\mathbb{R}^{3}\otimes(\mathfrak{g}^{\perp}_{m})^{*}.$
The equation $\mu(q)=-\zeta$ then becomes
$\hat{\mu}_{m}(q)=-\zeta_{m}\quad\mbox{and}\quad\check{\mu}_{m}(q)=-\zeta_{m}^{\perp},$
where $\zeta=(\zeta_{m},\zeta_{m}^{\perp})$ in the decomposition (4.16).
We will use this decomposition to see that $\widetilde{\mu^{-1}(-\zeta)}$ is a $p$-submanifold near $H_{i}$, but in order to do that, we need to introduce suitable coordinates. Let $\mathbb{S}(\mathbb{M})$ be the sphere of radius one centered at the origin in $\mathbb{M}$. Without loss of generality we can think of $m$ as a point on that sphere. Since the action of $G$ is unitatry, its action restricts to $\mathbb{S}(\mathbb{M})$. Applying the tube theorem [12, Theorem 2.4.1] near $m$ on $\mathbb{S}(\mathbb{M})$, we can find a local neighborhood $\mathcal{U}$ of $m$ in $\mathbb{S}(\mathbb{M})$ with an equivariant diffeomorphism
$\mathcal{U}\cong G\times_{G_{m}}V$
with $G_{m}$ the stabilizer of $m$, where $V$ is a linear representation of $G_{m}$. Let
$\mathcal{C}_{\mathcal{U}}:=\{\lambda p\;|\;p\in\mathcal{U},\;\lambda\geq 0\}$
be the corresponding cone in $\mathbb{M}$ with cross-section $\mathcal{U}$. Since $\mu$ is equivariant and $\zeta\in\mathbb{R}^{3}\otimes Z$ with $Z$ the center of $\mathfrak{g}$, to check that $\widetilde{\mu^{-1}(-\zeta)}\cap\mathcal{C}_{\mathcal{U}}$ is a $p$-submanifold near $H_{i}$, it sufffices to check that $\widetilde{\mu^{-1}(-\zeta)}\cap\mathcal{C}_{V}$ is a $p$-submanifold near $H_{i}$, where $\mathcal{C}_{V}$ is the cone over $V$ seen as the subset $\{e\}\times V$ of $\mathcal{U}$. Let $\check{V}$ be the subspace of $V$ on which $G_{m}$ acts trivially, so that there is a decomposition
$V=\check{V}\oplus\hat{V}$
with $\hat{V}$ a linear representation of $G_{m}$ such that the origin is the only point with stabilizer $G_{m}$. Hence, $\check{V}$ can be identified with the stratum $\overline{s}_{i}\cap\overline{\mathcal{C}}_{V}$ and there are natural identifications $\check{V}=\left(\bigoplus\limits_{i=1}^{3}I_{i}T_{m}\mathcal{O}_{m}\right)\oplus T_{m}$ and $\hat{V}=T_{m}^{\perp}$. Since $\hat{\mu}_{m}$ is the moment map for the action of $G_{m}$ and that $\mu(m)=0$, notice that $\hat{\mu}_{m}$ vanishes quadratically on $\check{V}$. If $\mu_{V}$ is the restriction of the moment map to $V$, then by homogeneity, on $C_{V}$, $\mu$ is given by
$\mu(\rho,\omega)=\rho^{2}\mu_{V}(\omega)$
for $\omega\in V$ and $\rho$ the distance function from the origin in $\mathbb{M}$. If $\check{\omega}$ and $\hat{\omega}$ are the coordinates on $\check{V}$ and $\hat{V}$ respectively, notice that the blow-up of $\overline{s}_{i}$ at infinity corresponds to introducing the coordinates $(u,\check{\omega},\hat{v})$ with $u=\rho^{-1}$ and $\hat{v}=\frac{\hat{\omega}}{u}$. In these coordinates, the fiber bundle $\phi_{i}:H_{i}\to S_{i}$ corresponds to the projection
$(\check{\omega},\hat{v})\mapsto\check{\omega}.$
Since $\hat{\mu}_{m}$ vanishes quadratically on $\check{V}$, notice that in terms of these coordinates,
$\hat{\mu}_{m}(u,\check{\omega},\hat{v})=\hat{\nu}_{0}(\check{\omega},\hat{v})+u\hat{f}(u,\check{\omega},\hat{v})$
for $\hat{\nu}_{0}$ a smooth function which is quadratic in $\hat{v}$ and $\hat{f}$ a smooth function. On the other hand, expanding $\check{\mu}_{m}$ in Taylor series at $\check{V}\times\{0\}\subset V$, we see that
$\check{\mu}_{m}(u,\check{\omega},\hat{v})=u^{-2}\check{\nu}_{2}(\check{\omega})+u^{-1}\check{\nu}_{1}(\check{\omega},\check{v})+\check{\nu}_{0}(u,\check{\omega},\hat{v}),$
where the $\check{\nu}_{i}$ are smooth functions with $\check{\nu}_{1}$ linear in $\hat{v}$. Thus, in these coordinates, $\mu^{-1}(\zeta)$ corresponds to the solutions of the equations
(4.17) $\check{\nu}_{2}(\check{\omega})+u\check{\nu}_{1}(\check{\omega},\check{v})+u^{2}\check{\nu}_{0}(u,\check{\omega},\hat{v})+u^{2}\zeta_{m}^{\perp}$ $=0,$
$\hat{\nu}_{0}(\check{\omega},\hat{v})+u\hat{f}(u,\check{\omega},\hat{v})$ $=-\zeta_{m}.$
These equations also make sense at $u=0$, in which case we obtain
(4.18) $\check{\nu}_{2}(\check{\omega})$ $=0,$
$\hat{\nu}_{0}(\check{\omega},\hat{v})$ $=-\zeta_{m},$
with the first equation defining $\Sigma_{i}$ inside $S_{i}$ and the second equation defining $\widetilde{\mu^{-1}(\zeta)}\cap\phi_{i}^{-1}(\sigma)$ above $\sigma\in\Sigma_{i}$. By definition of a moment map, notice that the differential of the restriction of $\check{\mu}_{m}$ to $\check{V}$ as rank $3\dim T_{m}\mathcal{O}_{m}$, confirming that $\check{\nu}_{2}$ has differential of maximal rank on the locus defined by (4.18), in agreement with the fact that $\Sigma_{i}$ is smooth. On the other hand, above $m$, we have in fact that $\widetilde{\mu^{-1}(-\zeta)}\cap\phi_{i}^{-1}(m)$ is given by the equation $\hat{\mu}_{m}(\hat{v})=-\zeta_{m}$, which by proper genericity of $\zeta$ defines a smooth quiver variety. Since initially, $m\in\Sigma_{i}$ was arbitrary, we see by induction on depth that we can assume that $\widetilde{\mu^{-1}(\zeta)}\cap\phi_{i}^{-1}(\sigma)$ is a $p$-submanifold inside $\phi_{i}^{-1}(\sigma)$ for any $\sigma\in\Sigma_{i}$ with $\hat{\nu}_{0}(\check{\omega},\hat{v})$ having differential of maximal rank in (4.18). The local description (4.17) then shows that $\widetilde{\mu^{-1}(-\zeta)}$ is a p-submanifold near $H_{i}$ with a natural fiber bundle
$H_{i}\cap\widetilde{\mu^{-1}(-\zeta)}\to\Sigma_{i}$
induced by $\phi_{i}$.
Along a non-minimal boundary hypersurface $H_{i}$, we can assume by induction that we already know that $\widetilde{\mu^{-1}(-\zeta)}$ is smooth near $H_{j}$ for $H_{j}<H_{i}$. Thus, it suffices to show that $\widetilde{\mu^{-1}(-\zeta)}$ is smooth near $\phi_{i}^{-1}(m)$ for $m$ in the interior of $\Sigma_{i}$, so that the same argument as before applies. Clearly then $\widetilde{\mu^{-1}(-\zeta)}$ is smooth and has the claimed iterated fibration structure making it a manifold with fibered corners. Moreover, the free action of $G$ on $\mu^{-1}(-\zeta)$ extends to a free smooth action on $\widetilde{\mu^{-1}(-\zeta)}$ which is compatible with the iterated fibration structure. ∎
Corollary 4.7.
For $\zeta\in\mathbb{R}^{3}\otimes Z$ properly generic, the Nakajima metric on the quiver variety $\mathfrak{M}_{\zeta}$ is an exact quasi-asymptotically conical metric with smooth expansion at infinity.
Proof.
By [9], the Euclidean metric on $\mathbb{M}$ can be seen as an exact $\operatorname{QAC}$ metric on $\widetilde{\mathbb{M}}$. Since the iterated fibration structure of $\widetilde{\mu^{-1}(-\zeta)}$ is induced by the one of $\widetilde{\mathbb{M}}$ through the inclusion $\widetilde{\mu^{-1}(-\zeta)}\to\widetilde{\mathbb{M}}$, the restriction of the Euclidean metric of $\mathbb{M}$ to $\widetilde{\mu^{-1}(-\zeta)}$ is automatically an exact $\operatorname{QAC}$ metric. Since $G$ acts freely on $\widetilde{\mu^{-1}(-\zeta)}$ in a way compatible with the metric and the iterated fibration structure, this metrics descends to induce an exact $\operatorname{QAC}$ metric on the quotient $\widetilde{\mu^{-1}(\zeta)}/G$ with smooth asymptotic expansion at infinity. ∎
5. Wedge $3$-Sasakian manifolds
In this section, we will briefly review the notion of $3$-Sasakian manifold and allow for possible singularities of wedge type. In this singular setting, we will extend and refine some vanishing results in cohomology of Galicki and Salamon [13] relying on the Weitzenböck formula of Semmelmann and Weingart [32]. This will allow us to show that the associated Hodge-deRham operator is automatically essentially self-adjoint and has some spectral gaps in certain degrees. These results will allow us in the next section to apply the results of [20] to the Hodge-deRham operators appearing in the study of the Hodge cohomology of quiver varieties.
Consider a Riemannian manifold $(\mathcal{S},g)$ with Levi-Civita connection $\nabla$. For $\xi$ a vector field on $\mathcal{S}$, let $\eta$ denote the $1$-form dual to $\xi$ and let $\Xi$ be the endomorphism of the tangent bundle defined by $\Xi(X)=\nabla_{X}\xi$. Notice that $\xi$ will be a Killing vector field if and only if $\Xi$ is skew-symmetric.
Definition 5.1.
The triple $(\mathcal{S},g,\xi)$ is a Sasakian manifold if $\xi$ is a Killing vector field of unit length and
$(\nabla_{X}\Xi)(Y)=\eta(Y)X-g(X,Y)\xi$
for all vector fields $X$ and $Y$. In this case, we say $g$ is a Sasakian metric.
Referring to [5] and references therein for more details, let us recall that one of the main features of a Sasakian manifold is that the associated cone metric
$dr^{2}+r^{2}g$
on +×$\mathbb{R}^{+}\times\mathcal{S}$ is Kähler. In particular, Sasakian manifolds are always odd dimensional. In terms of the complex structure $J$ of the Kähler cone, the Killing vector field $\xi$ is then given by $J(r\frac{\partial}{\partial r})$ when $\mathcal{S}$ is identified with the cross-section $\{1\}\times\mathcal{S}$ of the cone, while the Kähler form of the Kähler cone metric is given by
$\omega=\frac{1}{2}\partial\overline{\partial}r^{2}.$
When the Kähler cone is Ricci-flat, the Sasakian manifold is Einstein with positive Einstein constant equal to $\dim\mathcal{S}-1$. Requiring furthermore that the Kähler cone be hyperKähler yields the following structure on $\mathcal{S}$.
Definition 5.2.
A $3$-Sasakian manifold is a Riemannian manifold $(\mathcal{S},g)$ such that the cone metric $dr^{2}+r^{2}g$ on +×$\mathbb{R}^{+}\times\mathcal{S}$ is hyperKähler. Equivalently, a $3$-Sasakian manifold is a Riemannian manifold $(\mathcal{S},g)$ admitting three distinct Sasakian structures with Killing vector fields $\xi^{1},\xi^{2}$ and $\xi^{3}$ mutually orthogonal and such that
(5.1) $[\xi^{a},\xi^{b}]=\sum_{c=1}^{3}\epsilon^{abc}\xi^{c}\quad\mbox{for}\quad a,b,c\in\{1,2,3\}.$
By (5.1), the vector fields $\xi^{1},\xi^{2}$ and $\xi^{3}$ generate a Lie algebra naturally isomorphic to the Lie algebra of $\operatorname{Sp}(1)$. In fact, by Frobenius theorem, they induce a foliation $\mathcal{F}$ on $\mathcal{S}$ and correspond to the infinitesimal generators of an action of $\operatorname{Sp}(1)$ on $\mathcal{S}$ with the leaves of $\mathcal{F}$ corresponding to the orbits of this $\operatorname{Sp}(1)$-action. The $3$-Sasakian structure behaves nicely with respect to this foliation. More precisely, referring to [13, Theorem 1.3] or [5, Proposition 13.3.11 and Theorem 13.3.13] for further details and references, there is the following well-known result.
Theorem 5.3.
Let $(\mathcal{S},g,\xi^{a})$ be a $3$-Sasakian manifold of dimension $34n+$ 3 such that the vector fields $\xi^{1},\xi^{2}$ and $\xi^{3}$ are complete. Then:
  1. (1)
    $g$ is Einstein with scalar curvature $2(2n+1)(4n+3)$;
  2. (2)
    The foliation $\mathcal{F}$ generated by $\xi^{1},\xi^{2}$ and $\xi^{3}$ is Riemannian with respect to the metric $g$;
  3. (3)
    Each leaf is totally geodesic and of constant curvature $1$, while the space of leaves $Q$ is a quaternionic-Kähler orbifold of scalar curvature $16n(n+2)$;
  4. (4)
    The natural projection $\pi:\mathcal{S}\to Q$ is a principal orbibundle with group $\operatorname{Sp}(1)$ or $\operatorname{SO}(3)$.
In particular, a $3$-Sasakian manifold is Einstein with positive Einstein constant $\dim\mathcal{S}-1$. By Myers’s theorem, a complete $3$-Sasakian manifold is therefore compact and has finite fundamental group. This implies that its first Betti number vanishes. More generally, it was shown by Galicki and Salamon [13, Theorem A] that the odd Betti numbers $b_{2k+1}$ of $\mathcal{S}$ vanish for $0\leq k\leq\frac{\dim\mathcal{S}-3}{4}$.
Motivated by the study of quiver varieties, where the hyperKähler cones showing up are typically singular, we want to extend and refine this vanishing result to singular $3$-Sasakian metrics. We will concentrate our effort on the case where the singular metric is an exact wedge metric.
Definition 5.4.
An exact wedge metric $g_{w}$ on a manifold with fibered corners $\mathcal{S}$ is $3$-Sasakian if $(\mathcal{S}\setminus\partial\mathcal{S},g_{w})$ is $3$-Sasakian with Killing vector fields $\xi^{1},\xi^{2}$ and $\xi^{3}$ extending to complete wedge vector fields in $\mathcal{C}^{\infty}(\mathcal{S};{}^{w}T\mathcal{S})$ generating an action of $\operatorname{Sp}(1)$ such that for each $H\in\mathcal{M}_{1}(\mathcal{S})$:
  1. (1)
    If $a_{e}:{}^{e}T\mathcal{S}\to T\mathcal{S}$ denotes the anchor map of the edge cotangent bundle, then for $a\in\{1,2,3\}$, $a_{e}(v\xi^{a})|_{H}=0$ and $\xi^{a}$ descends to a wedge vector field $\xi^{a}_{S_{H}}\in\mathcal{C}^{\infty}(S_{H};{}^{w}TS_{H})$ on the base $S_{H}$ of $\phi_{H}:H\to S_{H}$ making the exact wedge metric $g_{S_{H},w}$ on $S_{H}$ induced by $g_{w}$ a $3$-Sasakian metric with Killing vector fields $\xi^{1}_{S_{H}},\xi^{2}_{S_{H}}$ and $\xi^{3}_{S_{H}}$;
  2. (2)
    For each $s\in S_{H}$, the exact wedge metric $g_{w,s}$ on the manifold with fibered corners $\phi_{H}^{-1}(s)$ induced by $g_{w}$ is such that
    $dx_{H}^{2}+x_{H}^{2}g_{w,s}$
    is a hyperKähler cone making $g_{w,s}$ an exact wedge $3$-Sasakian metric.
Remark 5.5.
Since the fibers and the base of $\phi_{H}:H\to S_{H}$ have depth lower than $\mathcal{S}$, the definition above is not circular, namely proceeding by induction on the depth of $\mathcal{S}$, we can assume that the notion of $3$-Sasakian exact wedge metric is well-defined on manifolds with fibered corners of lower depth.
Models at infinity of quiver varieties yield natural example of wedge exact $3$-Sasakian metrics.
Example 5.6.
Let $\mathfrak{M}_{\zeta}$ be a quiver variety as in Corollary 4.7. Then the model wedge exact metric $g_{w}$ in (3.9) for the maximal hypersurface of the $\operatorname{QAC}$ compactification of $\mathfrak{M}_{\zeta}$ is a wedge exact $3$-Sasakian metric. Indeed, by Corollary 4.7, the cone metric $dr^{2}+r^{2}g_{w}$ is hyperKähler since it corresponds to the singular quiver variety $\mathfrak{M}_{0}$. By the construction leading to Theorem 4.6 and Corollary 4.7, condition (2) of Definition 5.4 holds with the hyperKähler cone $\hat{\mu}^{-1}_{m}(0)/G_{m}$ for $m\in\Sigma_{i}$ representing a point in the base $\Sigma_{i}/G$ of a boundary hypersurface $H_{i}\cap\widetilde{\mu^{-1}(\zeta)}/G$ of $\widetilde{\mu^{-1}(\zeta)}/G$. On the other hand, condition (1) in Definition 5.4 follows from a result of Dancer-Swann [10] (see also [22, Theorem 1.1]) applied to the hyperKähler cone $\mathfrak{M}_{0}$. This shows at the same time that exact wedge metrics of (3.9) for the other boundary hypersurfaces of the $\operatorname{QAC}$ compactification of $\mathfrak{M}_{\zeta}$ are also exact wedge $3$-Sasakian manifolds.
Theorem 5.3 naturally extends to $3$-Sasakian exact wedge metrics. Indeed, by Definition 5.4, the vector fields $\xi^{1}$, $\xi^{2}$ and $\xi^{3}$ are are complete wedge vector fields on $\mathcal{S}$, i.e. they are also smooth vector fields on $\mathcal{S}$ and their flows exist for all time on $\mathcal{S}$, so generate a locally free action of $\operatorname{Sp}(1)$ on $\mathcal{S}$ such that for each $H\in\mathcal{M}_{1}(\mathcal{S})$, there is a corresponding action on $S_{H}$ making the map $\phi_{H}:H\to S_{H}$ $\operatorname{Sp}(1)$-equivariant. The $\operatorname{Sp}(1)$-action also descends to a $\operatorname{Sp}(1)$-action on the smoothly stratified space $\hat{\mathcal{S}}$ associated to $\mathcal{S}$. Since the action is locally free and $\operatorname{Sp}(1)$ is compact, the foliation induced by the orbits of this action is automatically quasi-regular, that is, the quotient $\mathcal{S}/\operatorname{Sp}(1)$ is a orbifold with corners in the sense of [9]. Since the action is compatible with the iterated fibration structure of $\mathcal{S}$, notice that $\hat{Q}=\hat{\mathcal{S}}/\operatorname{Sp}(1)$ is naturally a smoothly stratified space with strata coming from those of $\mathcal{S}$ and orbifold singularities created by taking the quotient of the $\operatorname{Sp}(1)$-action. More precisely, by a result of Tanno [34], see also [5, Proposition 13.3.11], if the $\operatorname{Sp}(1)$ action is nowhere free, then the smallest conjugacy class of the stabilizer groups is the one of $\mathbb{Z}_{2}\subset\operatorname{Sp}(1)$. Thus, setting
$G=\left\{\begin{array}[]{ll}\operatorname{SO}(3)=\operatorname{Sp}(1)/\mathbb{Z}_{2},&\mbox{the $\operatorname{Sp}(1)$-action is nowhere free},\\ \operatorname{Sp}(1),&\mbox{otherwise},\end{array}\right.$
the stratification on $\mathcal{S}$ induced by the conjugacy classes of stabilizer subgroups of the $\operatorname{Sp}(1)$-action is given by strata of the form
$s_{I}=\{p\in\mathcal{S}\;|\;G_{p}\in I\}$
for I a conjugacy class of subgroups in $G$ and $G_{p}\subset G$ the stabilizer group of $p$ in $G$. Let $\overline{s}_{I}$ be the closure $s_{I}$ in $\mathcal{S}$. To resolve the $G$-action on $\mathcal{S}$ into a free action, we could as in [2] blow up the closed strata $\overline{s}_{I}$ in an order compatible with the partial order on the strata given by
$s_{I}<s_{J}$ $\Longleftrightarrow\;s_{I}\subsetneq\overline{s}_{J}$
$\Longleftrightarrow\;\mbox{For $K\in J$, there exists $L\in I$ such that $K\subset L$.}$
However, to describe $\hat{\mathcal{S}}/\operatorname{Sp}(1)=\hat{\mathcal{S}}/G$ as a smoothly stratified space, each stratum $s_{I}$ needs to be decomposed as a disjoint union
$s_{I}=s_{I,\mathcal{S}}\sqcup\left(\bigsqcup_{H\in\mathcal{M}_{1}(\mathcal{S})}s_{I,H}\right),$
where
$s_{I,H}=s_{I}\cap\left(H\setminus\left(\bigcup_{L<H}L\cap H\right)\right)$
for $H\in\mathcal{M}_{1}(\mathcal{S})$ and where $s_{I,\mathcal{S}}=s_{I}\setminus(s_{I}\cap\partial\mathcal{S})$. Proceeding lexicographically, there is a partial order on this refined decomposition given by
(5.2) $s_{I,H}<s_{J,L}\;\Longleftrightarrow\;I<J,\quad\mbox{or $I=J$ and $H<L$},$
where we used the convention that $H<\mathcal{S}$ for all $H\in\mathcal{M}_{1}(\mathcal{S})$ when $L=\mathcal{S}$. If $\overline{s}_{I,H}$ denotes the closure of $s_{I,H}$, then we can resolve the $G$-action into a free action on the space
(5.3) $Y=[\mathcal{S};\{\overline{s}_{I,H}\},I\in\mathcal{I}\setminus\{I_{\operatorname{Id}}\},\;H\in\mathcal{M}_{1}(\mathcal{S})\cup\{\mathcal{S}\}],$
obtained from $\mathcal{S}$ by blowing up the $\overline{s}_{I,H}$ in an order compatible with the partial order described above, where $\mathcal{I}$ is the set of conjugacy classes of subgroups of $G$ and $I_{\operatorname{Id}}$ is the conjugacy class corresponding to the trivial subgroup $\{\operatorname{Id}\}$. One can readily check that the quotient
$Q=Y/G$
is naturally a manifold with fibered corners with associated smoothly stratified space $\hat{Q}=\hat{\mathcal{S}}/G$. This yields the following generalization of Theorem 5.3.
Theorem 5.7.
Let $g_{w}$ be a $3$-Sasakian exact wedge metric on a manifold with fibered corners $\mathcal{S}$ of dimension $34n+$ 3. Then:
  1. (1)
    $g_{w}$ is Einstein with scalar curvature $2(2n+1)(4n+3)$;
  2. (2)
    The foliation $\mathcal{F}$ generated by $\xi^{1},\xi^{2}$ and $\xi^{3}$ is Riemannian with respect to the metric $g_{w}$ on $\mathcal{S}\setminus\partial\mathcal{S}$ and with respect to $g_{S_{H},w}$ on $S_{H\setminus}\partial S_{H}$ for each $H\in\mathcal{M}_{1}(\mathcal{S})$;
  3. (3)
    Each leaf is totally geodesic and of constant curvature $1$, while on the quotient $Q=Y/G$, the metric $g_{w}$ induces a quaternionic-Kähler exact wedge metric of scalar curvature $16n(n+2)$;
  4. (4)
    The natural projection $\pi:\mathcal{S}\to\mathcal{S}/\operatorname{Sp}(1)$ is a principal orbibundle with group $\operatorname{Sp}(1)$ or $\operatorname{SO}(3)$.
Applying the Weitzenböck formula of Semmelmann and Weingart [32] to the quaternionic-Kähler wedge metric of Theorem 5.7, we obtain the following.
Theorem 5.8 ([32]).
Let $g_{w,QK}$ be the Quaternionic-Kähler exact wedge metric on $Q=Y/G$ of Theorem 5.7. Let $F\to Q$ be a flat Euclidean vector bundle and let ð$\eth_{w,QK}$ is the Hodge-deRham operator associated to $g_{w,QK}$ and $F$. Then, for $0\leq k\leq n$,
(5.4) $\langle\psi,\eth^{2}_{w,QK}\psi\rangle_{L^{2}_{w}}\geq 2\langle\psi,\psi\rangle_{L^{2}_{w}}\quad\forall\psi\in\Omega^{2k+1}_{c}(Q\setminus\partial Q;F),$
where $\langle\cdot,\cdot\rangle_{L^{2}_{w}}$ is the $L^{2}$-inner product associated to $g_{w,QK}$ and the bundle metric of $F$ and $\Omega^{q}_{c}(Q\setminus\partial Q;F)$ is the space of compactly supported smooth $F$-valued forms on $Q\setminus\partial Q$.
Proof.
In [32], Semmelmann and Weingart give a detailed description of the curvature term $R_{QK}$ in the Weitzenböck formula
(5.5) ð$\eth_{w,QK}^{2}=\nabla^{*}\nabla+R_{QK}$
by decomposing it in terms of the irreducible representations of the holonomy group $\operatorname{Sp}(1)\cdot\operatorname{Sp}(n)$ of $g_{w,QK}$. No flat Euclidean vector bundle was considered in [32], but since the formula is local, notice that it also holds for the Hodge-deRham operator acting of $F$-valued forms. For $0\leq k\leq n$, they obtain the following estimate on $R_{QK}$ acting on ($F$-valued) forms of degree $12k+$ 1,
(5.6) $R_{QK}\geq\frac{\kappa_{w,QK}}{8n(n+2)},$
where $\kappa_{w,QK}$ is the scalar curvature of $g_{w,QK}$. This estimate is not explicitly written in [32], but it follows from [32, Lemma 5.2] combined with [32, (22)], [32, Proposition 5.1] and the way [32, Theorem 3.4] is used in its proof. Since $\kappa_{w,QK}=16n(n+2)$ by Theorem 5.7, this means that
(5.7) $R_{QK}\geq 2.$
The result is then a direct consequence of (5.5) and (5.7).
Remark 5.9.
In [32, Proposition 5.4], the estimate (5.6) is used to show that a closed Quaternionic-Kähler manifold of dimension $4n$ has no deRham cohomology in degree $12k+$ 1 for $k\in\{0,\ldots,n\}$.
The previous result can be lifted to a $3$-Sasakian manifold as follows.
Theorem 5.10.
Let $g_{w}$ be an exact wedge $3$-Sasakian metric on a manifold with fibered corners $\mathcal{S}$ of dimension $34n+$ 3. Let $F\to\mathcal{S}$ be a flat Euclidean vector bundle and let ð$\eth_{w}$ be the Hodge-deRham operator associated to $g_{w}$ and $F$. Then for $k\in\{0,\ldots,n\}$,
(5.8) ð$\langle\eth_{w}\psi,\eth_{w}\psi\rangle_{L^{2}_{w}}\geq 2\langle\psi,\psi\rangle_{L^{2}_{w}}\quad\forall\psi\in\Omega^{2k+1}_{c}(\mathcal{S}\setminus\partial\mathcal{S};F).$
In particular, when $\mathcal{S}$ is a closed manifold, this implies the vanishing theorem of Galicki-Salamon, namely that the space of harmonic $F$-valued forms in degree $12k+$ 1 is trivial for $k\in\{0,\ldots,n\}$.
Proof.
If $\psi$ is of vertical degree $1$ with respect to the foliation $\mathcal{F}$, so of horizontal degree $2k+1-1=2k$, then if $\eth_{\mathcal{F}}$ denotes the vertical part of the Hodge-deRham operator, namely the part that only changes the vertical degree, then
(5.9) ð$\langle\eth_{w}\psi,\eth_{w}\psi\rangle_{L^{2}_{w}}\geq\langle\eth_{\mathcal{F}}\psi,\eth_{\mathcal{F}}\psi\rangle_{L^{2}_{w}}.$
However, combining the Weitzenböck formula on $1$-forms, which also holds when the $1$-forms are $F$-valued, together with the fact that the leaves of the foliation $\mathcal{F}$ are of constant curvature $1$, so Einstein with Einstein constant $2$, we have that
(5.10) ð$\langle\eth_{\mathcal{F}}\psi,\eth_{\mathcal{F}}\psi\rangle_{L^{2}_{w}}$ $\geq\langle\psi,\operatorname{Ric}_{\mathcal{F}}\psi\rangle_{L^{2}_{w}}$
$\geq 2\langle\psi,\psi\rangle_{L^{2}_{w}},$
where $\operatorname{Ric}_{\mathcal{F}}$ is the Ricci curvature form of the leaves of $\mathcal{F}$. Hence, the result follows in this case from (5.9) and (5.10). If instead $\psi$ is of vertical degree $2$, then using the symmetry induced by the Hodge star operator in the leaves, we see that the same argument can be used to obtain
(5.11) ð$\langle\eth_{w}\psi,\eth_{w}\psi\rangle_{L^{2}_{w}}\geq\langle\eth_{\mathcal{F}}\psi,\eth_{\mathcal{F}}\psi\rangle_{L^{2}_{w}}\geq 2\langle\psi,\psi\rangle_{L^{2}_{w}}.$
If $\psi$ is of vertical degree 00, so of horizontal degree $12k+$ 1, we need to proceed differently. Notice first that on a given leaf $\ell$ of $\mathcal{F}$,
$\ell\cong\operatorname{Sp}(1)/\Gamma$
for some finite subgroup $\Gamma\subset\operatorname{Sp}(1)$ and that
$F|_{\ell}\cong\operatorname{Sp}(1)\times_{\Gamma}V$
for some orthogonal representation $V$ of $\Gamma$. In particular, the $\operatorname{Sp}(1)$-action (on the left of $\ell$) naturally extends to an action on $F|_{\ell}$. Clearly, these actions combine to give a global $\operatorname{Sp}(1)$-action on $F$ making the bundle projection $F\to\mathcal{S}$ $\operatorname{Sp}(1)$-equivariant. This induces a natural $\operatorname{Sp}(1)$-action on $\Omega_{c}^{*}(\mathcal{S};F)$ and we can consider the orthogonal decomposition of $\psi$ given by
$\psi=\psi_{1}+\psi_{2}$
with $\psi_{1}$ the $\operatorname{Sp}(1)$-invariant part of $\psi$. On the universal cover ~$\widetilde{\ell}\cong\operatorname{Sp}(1)$ of a leaf $\ell$, the pull-back $\psi_{2,\widetilde{\ell}}$ of $\psi_{2}|_{\ell}$ to $\widetilde{\ell}$ is orthogonal to the $\operatorname{Sp}(1)$-invariant horizontal forms of the degree $12k+$ 1, so orthogonal to the kernel of the lift of $\eth_{\mathcal{F}}|_{\ell}$ to $\widetilde{\ell}$. Using the fact that positive eigenvalues of ð$\eth_{\mathcal{F}}^{2}$ in vertical degree 00 correspond to a subset of the eigenvalues of $\eth_{\mathcal{F}}|_{\ell}$ on forms of vertical degree $1$, we deduce from (5.10) that
(5.12) ð$\langle\eth_{w}\psi_{2},\eth_{w}\psi_{2}\rangle_{L^{2}_{w}}\geq\langle\eth_{\mathcal{F}}\psi_{2},\eth_{\mathcal{F}}\psi_{2}\rangle_{L^{2}_{w}}\geq 2\langle\psi_{2},\psi_{2}\rangle_{L^{2}_{w}}.$
For the $\operatorname{Sp}(1)$-invariant part $\psi_{1}$, it corresponds to a $(2k+1)$-form on $Q$ taking values in a flat Euclidean vector orbibundle. Since orbifold singularities do not compromise the essential self-adjointness of the Hodge-deRham operator, we can without loss of generality suppose that $\psi_{1}$ is supported away from the orbifold singularities, so by Theorem 5.8, we obtain that
(5.13) ð$\langle\eth_{w}\psi_{1},\eth_{w}\psi_{1}\rangle_{L^{2}_{w}}\geq 2\langle\psi_{1},\psi_{1}\rangle_{L^{2}_{w}}.$
The result when $\psi$ is of vertical degree zero then follows from (5.12) and (5.13).
Finally, if $\psi$ is of vertical degree $3$, so of horizontal degree $22k-$ 2, then we can again decompose $\psi$ orthogonally, namely
$\psi=\psi_{1}+\psi_{2}$
with $\psi_{1}$ the $\operatorname{Sp}(1)$-invariant part of $\psi$. For $\psi_{2}$, its lift to the universal cover $\widetilde{\ell}$ of a leaf $\ell$ of $\mathcal{F}$ is again orthogonal to the kernel of the lift of $\eth_{\mathcal{F}}|_{\ell}$ to $\widetilde{\ell}$. Now, the positive eigenvalues of the lift of ð$\eth_{\mathcal{F}}^{2}|_{\ell}$ to $\widetilde{\ell}$ in vertical degree $3$ correspond to a subset of the eigenvalues in degree $2$, so using (5.11) for forms of vertical degree $2$, we deduce that
(5.14) ð$\langle\eth_{w}\psi_{2},\eth_{w}\psi_{2}\rangle_{L^{2}_{w}}\geq\langle\eth_{\mathcal{F}}\psi_{2},\eth_{\mathcal{F}}\psi_{2}\rangle_{L^{2}_{w}}\geq 2\langle\psi_{2},\psi_{2}\rangle_{L^{2}_{w}}.$
For $\psi_{1}$, it must be of the form
$\psi_{1}=\Upsilon\wedge\hat{\psi}_{1},$
where $\hat{\psi}_{1}$ a $\operatorname{Sp}(1)$-invariant horizontal form of degree $22k-$ 2 and
$\Upsilon:=\eta^{1}\wedge\eta^{2}\wedge\eta^{3}$
is a $\operatorname{Sp}(1)$-invariant form of vertical degree $3$, where $\eta^{a}$ is the $1$-form dual to $\xi^{a}$. As shown in [13, (10)], the exterior derivative of $\Upsilon$ is given by
$d\Upsilon=\eta^{1}\wedge\eta^{2}\wedge\overline{\phi}^{3}+\eta^{2}\wedge\eta^{3}\wedge\overline{\phi}^{1}+\eta^{3}\wedge\eta^{1}\wedge\overline{\phi}^{2},$
where
$\overline{\phi}^{a}=d\eta^{a}+\sum_{b,c}\epsilon^{abc}\eta^{b}\wedge\eta^{c}$
is a $2$-form which is horizontal with respect to the foliation $\mathcal{F}$. Since the exterior product of a basic form of degree $22k-2\leq 2n-$ 2 by $\overline{\phi}^{a}$ can only increase the $L^{2}$-norm, we see that
(5.15) ð$\langle\eth_{w}\psi_{1},\eth_{w}\psi_{1}\rangle_{L^{2}_{w}}$ $\geq\langle d\psi_{1},d\psi_{1}\rangle_{L^{2}_{w}}\geq\langle(d\Upsilon)\wedge\hat{\psi}_{1},(d\Upsilon)\wedge\hat{\psi}_{1}\rangle_{L^{2}_{w}}$
$\geq 3\langle\hat{\psi}_{1},\hat{\psi}_{1}\rangle_{L^{2}_{w}}=3\langle\psi_{1},\psi_{1}\rangle_{L^{2}_{w}}$
$\geq 2\langle\psi_{1},\psi_{1}\rangle_{L^{2}_{w}}.$
The result when $\psi$ is of vertical degree $3$ therefore follows from (5.14) and (5.15). ∎
This result allows us to obtain nice mapping properties for the Hodge-deRham operator. To describe them, it is convenient as in [20, § 8] to introduce the conjugated operator
$D_{w}:=\left({\prod_{H\in\mathcal{M}_{1}(\mathcal{S})}x_{H}^{\frac{f_{H}+1}{2}}}\right)\eth_{w}\left({\prod_{H\in\mathcal{M}_{1}(\mathcal{S})}x_{H}^{-\frac{f_{H}+1}{2}}}\right),$
where $f_{H}$ is the dimension of the fibers of $\phi_{H}:H\to S_{H}$. The choice of weight is to ensure that $D_{w}$ acting formally on
$L^{2}_{b}(\mathcal{S};\Lambda^{*}({}^{w}T^{*}\mathcal{S})\otimes F)=\left({\prod_{H\in\mathcal{M}_{1}(\mathcal{S})}x_{H}^{\frac{f_{H}+1}{2}}}\right)L^{2}_{w}(\mathcal{S};\Lambda^{*}({}^{w}T^{*}\mathcal{S})\otimes F)$
corresponds to ð$\eth_{w}$ acting formally on $L^{2}_{w}(\mathcal{S};\Lambda^{*}({}^{w}T^{*}\mathcal{S})\otimes F)$. If we consider the cone metric $dr^{2}+r^{2}g_{w}$ on $\mathcal{C}=[0,\infty)_{r}\times\mathcal{S}$, we can consider the corresponding Hodge-deRham operator ð$\eth_{\mathcal{C}}$ acting on $F$-valued forms and its conjugated operator
$D_{\mathcal{C}}=\left({r^{\frac{\dim\mathcal{S}+1}{2}}\prod_{H\in\mathcal{M}_{1}(\mathcal{S})}x_{H}^{\frac{f_{H}+1}{2}}}\right)\eth_{\mathcal{C}}\left({r^{-\frac{\dim\mathcal{S}+1}{2}}\prod_{H\in\mathcal{M}_{1}(\mathcal{S})}x_{H}^{\frac{f_{H}+1}{2}}}\right).$
To $D_{\mathcal{C}}$, we can associate an indicial family
(5.16) $I(rD_{C},\lambda):=r^{-\lambda}(rD_{\mathcal{C}})r^{\lambda}.$
If $\overline{\eta}$ is a $F$-valued $k$-form on the cone $\mathcal{C}$ obtained by parallel transport of its restriction $\eta$ to $\{1\}\times\mathcal{S}$ along geodesics emanating from the apex of the cone, then there is a decomposition
$\overline{\eta}=\overline{\alpha}+dr\wedge\overline{\beta},\quad\eta=\alpha+dr\wedge\beta,\quad\overline{\alpha}=r^{k}\alpha,\quad\overline{\beta}=r^{k-1}\beta$
for some $F$-valued forms $\alpha$ and $\beta$ on $\mathcal{S}$. As described in [20, § 8], in terms of such a decomposition, the indicial family (5.16) takes the form
$I(rD_{\mathcal{C}},\lambda)=\left({\begin{array}[]{cc}D_{w}&-\lambda+N_{w}-\frac{\dim\mathcal{S}-1}{2}\\ \lambda+N_{w}-\frac{\dim\mathcal{S}+1}{2}&-D_{w}\end{array}}\right),$
where $N_{w}$ is the degree operator multiplying a form (of pure degree) on $\mathcal{S}$ by its degree. If we compose by Clifford multiplication by $dr$, that is, by
$c:=c(dr)=\left({\begin{array}[]{cc}0&-1\\ 1&0\end{array}}\right),$
this induces a formally self-adjoint operator
$cI(rD_{\mathcal{C}},\lambda)$
acting on $L^{2}_{b}(\mathcal{S};\Lambda^{*}({}^{r}T^{*}\mathcal{S})\otimes F)$ for $\lambda\in\mathbb{R}$. If we suppose that $cI(rD_{\mathcal{C}},0)$ is esentially self-adjoint, an indicial root of $I(rD_{\mathcal{C}},\lambda)$ is a value of $\lambda$ such that $cI(rD_{\mathcal{C}},\lambda)$ is not invertible as an operator
$cI(rD_{\mathcal{C}},\lambda):v^{\frac{1}{2}}\left({\prod_{H\in\mathcal{M}_{1}(\mathcal{S})}x_{H}^{\frac{f_{H}+1}{2}}}\right)H^{1}_{w}(\mathcal{S};\Lambda^{*}({}^{w}T^{*}\mathcal{S})\otimes F)\to v^{-\frac{1}{2}}L^{2}_{b}(\mathcal{S};\Lambda^{*}({}^{w}T^{*}\mathcal{S})\otimes F),$
where $v=\prod_{H\in\mathcal{M}_{1}(\mathcal{S})}x_{H}$.
Theorem 5.11.
Let $\mathcal{S}$ be a manifold with fibered corners of dimension $34n+$ 3 and $F\to\mathcal{S}$ a flat Euclidean vector bundle. Suppose that $g_{w}$ is an exact wedge $3$-Sasakian metric on $\mathcal{S}$ and let ð$\eth_{w}$ be the Hodge-deRham operator associated to $g_{w}$ and $F$. Then the wedge operator ð$\eth_{w}$ is essentially self-adjoint with self-adjoint extension $vH^{1}_{w}(\mathcal{S};\Lambda^{*}({}^{w}T^{*}\mathcal{S})\otimes F)$ and discrete spectrum. Moreover, the associated indicial family $I(rD_{\mathcal{C}},\lambda)$ is essentially self-adjoint for $\lambda=0$ and has no indicial root in $[0,1]$.
Proof.
By [20, Lemma 8.11], when $cI(rD_{\mathcal{C}},0)$ is essentially self-adjoint, the indicial roots of $I(rD_{\mathcal{C}},\lambda)$ are given by
(5.17) $\begin{gathered}\left\{q-\frac{\dim\mathcal{S}-1}{2},-\left({q-\frac{\dim\mathcal{S}+1}{2}}\right)\;|\;\ker_{L^{2}_{w}}\eth_{w}\;\mbox{is not trivial in degree $q$}\right\}\\ \bigcup\left\{\ell\pm\sqrt{\zeta+\left({q-\frac{\dim\mathcal{S}-1}{2}}\right)^{2}}\;|\;\ell\in\{0,1\},\;\zeta\in\operatorname{Spec}(\delta_{w}d)_{q}\setminus\{0\}\right\}\\ \bigcup\left\{\ell\pm\sqrt{\zeta+\left({q-\frac{\dim\mathcal{S}+1}{2}}\right)^{2}}\;|\;\ell\in\{0,1\},\;\zeta\in\operatorname{Spec}(d\delta_{w})_{q}\setminus\{0\}\right\},\\ \end{gathered}$
where $\delta_{w}$ is the formal adjoint of $d$ with respect to $g_{w}$ and $F$, while $(\delta_{w}d)_{q}$ and $(d\delta_{w})_{q}$ denote the parts of $\delta_{w}d$ and $d\delta_{w}$ acting on forms of degree $q$ (using the Friedrich extension). Now, if $\mathcal{S}$ is a closed manifold, then ð$\eth_{w}$ is automatically essentially self-adjoint. Moreover, by (5.17) and Theorem 5.10, the indicial family has no indicial root in $[0,1]$. More precisely, by (5.17), this amounts to show that $\ker_{L^{2}_{w}}\eth_{w}$ is trivial in degree $12n+$ 1 and $22n+$ 2 with $\inf\operatorname{Spec}(\eth^{2}_{w})>1$ in these degrees, a result which follows from Theorem 5.10 with $k=n$ in degree $12n+$ 1 and by symmetry using the Hodge star operator in degree $22n+$ 2.
Proceeding by induction on the depth, we can therefore assume that the theorem holds for manifolds of depth lower than the one of $\mathcal{S}$. In particular, we can assume that the result holds for the fibers of $\phi_{H}:H\to S_{H}$ for each $H\in\mathcal{M}_{1}(\mathcal{S})$ with the induced metric. This implies that [20, Assumption 8.9] holds for ð$\eth_{w}$ on $\mathcal{S}$, so [20, Corollaries 8.17 and 8.18] can be applied to conclude that ð$\eth_{w}$ is essentially self-adjoint with self-adjoint extension $vH^{1}_{w}(\mathcal{S};\Lambda^{*}({}^{w}T^{*}\mathcal{S})\otimes F)$ and discrete spectrum. This shows also that $cI(rD_{\mathcal{C}},0)$ is essentially self-adjoint. Using that ð$\eth_{w}$ is essentially self-adjoint, (5.17) and Theorem 5.10, we can conclude as before that $I(rD_{\mathcal{C}},\lambda)$ has no indicial root in $[0,1]$, completing the inductive step and the proof.
Remark 5.12.
To show essential self-adjointness and the discreteness of the spectrum, we can also use [1].
This yields the following generalization of the vanishing theorem of Galicki and Salamon [13].
Corollary 5.13.
Let $\mathcal{S}$, $g_{w}$ and $F$ be as in the previous theorem. Then for $k\in\{0,\ldots,n\}$, $\ker_{L^{2}_{w}}\eth_{w}$ is trivial in degree $12k+$ 1. By a result of Cheeger [1, Theorem 3.4], when $F$ is a trivial flat real line bundle, this implies that the lower and upper middle perversity intersection cohomology groups associated to the smoothly stratified space $\hat{\mathcal{S}}$ vanish in degree $12k+$ 1 for $k\in\{0,\ldots,n\}$.
6. Reduced $L^{2}$-cohomology of quiver varieties
The results of § 4 and § 5 ensure that we can use the pseudodifferential calculus of [20] to obtain the following result.
Theorem 6.1.
For $\zeta\in\mathbb{R}^{3}\times Z$ properly generic, the (possibly reduced) quiver variety $\mathfrak{M}_{\zeta}$ admits a $\operatorname{QAC}$ metric $\widetilde{g}_{\operatorname{QAC}}$ which is $\operatorname{QAC}$ equivalent to the Nakajima metric and such that its space of $L^{2}$-harmonic forms is finite dimensional and contained in $v^{\epsilon}L^{2}\Omega^{*}(\mathfrak{M}_{\zeta},\widetilde{g}_{\operatorname{QAC}})$ for some $\epsilon>0$. In particular, the reduced $L^{2}$-cohomology of the quiver variety is finite dimensional.
Proof.
By Corollary 4.7, the result will follow from [20, Theorem 17.5] provided that we can check that [20, Assumptions 17.1,17.2 and 17.4] hold for the hyperKähler metric on $\mathfrak{M}_{\zeta}$. When the $\operatorname{QAC}$ compactification $\widetilde{\mathfrak{M}_{\zeta}}=\widetilde{\mu^{-1}(\zeta)}/G$ of $\mathfrak{M}_{\zeta}$ is of depth 1, notice that [20, Assumption 17.1] is trivially satisfied. Proceeding by induction on the depth of $\widetilde{\mathfrak{M}_{\zeta}}$, we can suppose more generally that Theorem 6.1 holds for quiver varieties having a $\operatorname{QAC}$ compactification of lower depth. For $H\in\mathcal{M}_{1}(\widetilde{\mathfrak{M}_{\zeta}})$, this means that the hyperKähler metrics of the fibers of $\phi_{H}:H\to S_{H}$ have finite dimensional spaces of reduced $L^{2}$-cohomology, so finite dimensional spaces of $L^{2}$-harmonic forms. In particular, the dimension of these spaces only depend on the quasi-isometric class of the hyperKähler metric, hence, by the Ehresmann lemma of [19, Corollary A.6], this dimension is locally constant on $S_{H}$, showing that the fiberwise $L^{2}$-harmonic forms do form a finite rank vector bundle on $S_{H}$. This shows that [20, Assumption 17.1] holds in this case.
To complete the proof and the induction, we need to check that [20, Assumptions 17.2 and 17.4] also holds. First notice that by the result of Hitchin [16], the space of $L^{2}$-harmonic forms of a quiver variety is trivial except possibly in middle degree. In particular, all the spaces of $L^{2}$-harmonic forms occurring in [20, Assumptions 17.2 and 17.4] are trivial outside middle degree. In this case, one can check that [20, Assumptions 17.4] is implied by [20, Assumption 17.2], so we only need to check the latter. This assumption requires that some flat Euclidean vector bundles $F$ on some manifolds with fibered corners $\mathcal{S}$ have trivial spaces of $L^{2}$-harmonic forms in degree $q$ for
$\left|q-\frac{\dim\mathcal{S}}{2}\right|\leq 1$
with respect to some wedge metrics $g_{w}$. By Example 5.6, all these metrics are $3$-Sasakian exact wedge metrics, so $\dim\mathcal{S}$ is always odd and we need to check that the space of harmonic forms is trivial in degree $\frac{\dim\mathcal{S}\pm 1}{2}$. Now, by the symmetry of the Hodge star operator, we only need to check this in degree $\frac{\dim\mathcal{S}-1}{2}$, in which case the result follows from Corollary 5.13.
Remark 6.2.
Despite the spectral gap obtained in Theorem 5.10, we were not able to confirm that the decay of $L^{2}$-harmonic forms could directly apply to the hyperKähler metric of the quiver variety. One problem is that this spectral gap is not good enough to ensure that there are no indicial roots in the range prescribed by [18, Assumption 9.14] with $\delta=-\frac{1}{2}$.
Using the results of [18], this yields the following characterization of the reduced $L^{2}$-cohomology of a quiver variety.
Theorem 6.3.
If $(\mathfrak{M}_{\zeta},g_{N})$ is a (possibly reduced) quiver variety equipped with the Nakajima metric $g_{N}$ and with $\zeta$ properly generic, then
$\operatorname{Im}[H^{*}_{c}(\mathfrak{M}_{\zeta})\to H^{*}(\mathfrak{M}_{\zeta})]=\mathcal{H}^{*}(\mathfrak{M}_{\zeta}).$
Proof.
Since smooth quiver varieties are diffeomorphic to smooth affine complex varieties, they have no cohomology above middle degree by a result of Lefschetz [28, Theorem 7.2]. Using this property and Corollary 4.7, the proof of [18, Corollary 3.3] generalizes automatically quiver varieties with properly generic $\zeta$. Combined with Theorem 6.1, this allows to generalize the proof of [18, Theorem 3.5] to any quiver variety with properly generic $\zeta$, which yields the result.